### v1

# Diffusion Beats Autoregressive in Data-Constrained Settings

Abstract
Autoregressive (AR) models have long dominated the landscape of large language models, driving progress across a wide range of tasks. Recently, diffusion-based language models have emerged as a promising alternative, though their advantages over AR models remain underexplored. In this paper, we systematically study masked diffusion models in data-constrained settings—where training involves repeated passes over limited data—and find that they significantly outperform AR models when compute is abundant but data is scarce. Diffusion models make better use of repeated data, achieving lower validation loss and superior downstream performance. We interpret this advantage as implicit data augmentation: masked diffusion exposes the model to a diverse distribution of token orderings and prediction tasks, unlike AR’s fixed left-to-right factorization. We find new scaling laws for diffusion models and derive a closed-form expression for the critical compute threshold at which diffusion begins to outperform AR. These results suggest that when data, not compute, is the bottleneck, diffusion models offer a compelling alternative to the standard AR paradigm. Our code is available at: https://diffusion-scaling.github.io.
Figure 1: Pareto frontier of validation loss versus training FLOPs for autoregressive (AR) and masked diffusion models under data-constrained settings. Each point represents a model trained until convergence; we report the best validation loss achieved among all models using less than or equal to the FLOPs shown on the x-axis. AR models initially outperform diffusion models, particularly near the Chinchilla-optimal compute point (hoffmann2022chinchilla, ) (indicated on the plot). However, as training continues beyond this regime with repeated data, AR models quickly saturate and begin to overfit. In contrast, diffusion models continue to improve with more compute and exhibit no signs of overfitting.
1 Introduction
Training large language models (LLMs) on massive corpora of internet text has become the driver of recent AI breakthroughs (brown2020language, ; radford2019language, ; touvron2023llama, ). This progress has been fueled by scaling two core resources proportionately: compute and data  (kaplan2020scaling, ; hoffmann2022training, ). While compute availability has steadily grown—enabled by advances in hardware and the construction of larger data centers—the growth in high-quality data has been comparatively stagnant. Recent projections, such as those by Villalobos et.al. (villalobos2022will, ), estimate that the global supply of publicly available, human-generated data may be exhausted by 2028, posing a serious bottleneck to further scaling. This looming constraint makes it increasingly important to develop modeling strategies that are more data-efficient. Furthermore, there are several domains, such as robotics and healthcare, where the data, not compute, is a scarce resource even to begin with.
LLM development has so far been dominated by autoregressive (AR) models, which factorize the joint distribution of text in a fixed left-to-right order, i.e., $p(x)=p(x_{0})p(x_{1}\mid x_{0})p(x_{2}\mid x_{1})p(x_{3}\mid x_{1},x_{2})$. While this modeling approach has delivered state-of-the-art performance across a range of benchmarks, it remains unclear whether it is the optimal strategy going forward. Recently, diffusion-based models—specifically masked diffusion models  austin2021structured ; sahoo2024simple ; lou2310discrete ; shi2024simplified ; arriola2025block —have emerged as an alternative strategy, where they model text via a random-order factorization, e.g., $p(x)=p(x_{2}|x_{3})p(x_{3}|x_{0})p(x_{0}|x_{1})p(x_{1})$. These models learn across many different masked versions of the text, allowing them to train on diverse token orderings. Although diffusion language models have demonstrated similar scaling behavior to AR models (nie2024scaling, ; swerdlow2025unidisc, ), their practical benefits have, so far, been modest—largely due to their high training compute requirements.
This high compute demand has become the central obstacle to wider adoption of diffusion-based language models. As noted by Nie et al. (nie2024scaling, ) and Swerdlow et al. (swerdlow2025unidisc, ), masked diffusion models require up to 16× more compute than AR models to match validation NLL—a clear disadvantage for most applications.
But a critical nuance is often overlooked: these comparisons are based entirely on single-epoch training, where each token is seen only once. This conflates compute efficiency with data efficiency, making it unclear whether diffusion models truly need 16× more compute—or simply 16× more data.
To resolve this ambiguity, we systematically study masked diffusion models in data-constrained settings, where repeated training on limited data is the norm rather than the exception. We find that under such regimes, diffusion models substantially outperform autoregressive models across a variety of data scales and compute budgets. We train hundreds of models spanning multiple orders of magnitude in model size, data quantity, and number of training epochs to fit scaling laws for diffusion models in the data-constrained setting. We summarize some of our key findings below.
  1. 1.
    Diffusion models surpass autoregressive models given sufficient compute. Across a wide range of unique token budgets, we observe a consistent trend: autoregressive models initially outperform diffusion models at low compute, but quickly saturate. Beyond a critical compute threshold, diffusion models continue improving and ultimately achieve better performance (Section 4.1)
  2. 2.
    Diffusion models benefit far more from repeated data. Prior work (muennighoff2023scaling, ) showed that repeating the dataset up to 4 epochs is nearly as effective as using fresh data for autoregressive models. In contrast, we find that diffusion models can be trained on repeated data for up to 100 epochs, while having repeated data almost as effective as fresh data (Section 4.2).
  3. 3.
    Diffusion models have a much higher effective epoch count. Muennighoff et al. muennighoff2023scaling fit scaling laws for AR models in data-constrainted settings and define $R_{D}^{*}$ as a learned constant that characterizes the number of epochs after which training more epochs results in significantly diminished returns. For autoregressive models, they estimate $R_{D}^{*}\approx 15$ In contrast, we find $R_{D}^{*}\approx 500$ for diffusion models, suggesting they can benefit from repeated data over far more epochs without major degradation (Section 4.2).
  4. 4.
    Critical compute point follows a power law with dataset size. We find that the amount of compute required for diffusion models to outperform autoregressive models—the critical compute point—scales as a power law with the number of unique tokens. This yields a closed-form expression that predicts when diffusion becomes the favorable modeling choice for any given dataset size (Section 4.3).
  5. 5.
    Diffusion models yield better downstream performance. We find the above benefits extend beyond validation loss: the best diffusion model trained in data-constrained settings consistently outperform the best autoregressive model on a range of downstream language tasks (Section 4.4).
Through detailed scaling law analysis and downstream task evaluations, we demonstrate that diffusion models make significantly better use of repeated data, achieving lower validation loss and better generalization to downstream tasks. These results suggest that diffusion models may offer a compelling and underappreciated advantage in scenarios where data—not compute—is the primary bottleneck.
2 Related Work
Deep Learning in Data-Constrainted Settings.
Deep learning progress has been largely driven by the scaling of both data and compute. However, recent analyses suggest we may soon face a data bottleneck that could inhibit continued advancement (villalobos2022will, ). In language modeling, the dominant paradigm has been autoregressive (AR) models (vaswani2017attention, ; radford2019language, ; brown2020language, ), which are typically trained for a single epoch to maximize exposure to unique tokens (hoffmann2022training, ). In light of looming data constraints, Muennighoff et al.(muennighoff2023scaling, ) show that AR models can still benefit from data reuse: training for up to four epochs on repeated data achieves performance nearly on par with training on fresh data, suggesting an effective strategy for improving data efficiency. In contrast, computer vision has long embraced multi-epoch training along with aggressive data augmentation—such as random cropping, flipping, and color jittering—to expand effective dataset size and improve generalization(shorten2019survey, ; yang2022image, ), particularly for discriminative tasks like classification and detection. Despite these practices, data efficiency in generative modeling remains underexplored, and the trade-offs between leading paradigms such as diffusion and AR models under constrained data regimes are still poorly understood.
Diffusion-Based Language Models.
Diffusion models, originally developed for image generation (ho2020denoising, ), have recently been adapted to text, offering a fundamentally different paradigm for language modeling (austin2021structured, ; li2022diffusion, ; gulrajani2023likelihood, ). Broadly, diffusion language models fall into two categories: continuous and discrete. Continuous approaches (gulrajani2023likelihood, ) inject Gaussian noise in the forward process, whereas discrete methods (austin2021structured, ) corrupt tokens with noise sampled from distributions such as Bernoulli. Among the two classes, continuous diffusion has proven more difficult to scale on language data gulrajani2023likelihood ; lou2023discrete . In contrast, recent advances in discrete diffusion—particularly masked diffusion—have shown encouraging results. Recent work (arriola2025block, ; dieleman2023language, ; sahoo2024simple, ; lou2023discrete, ) has significantly narrowed the performance gap between diffusion and AR models. Notably, LLaDA (nie2025large, ) scales masked diffusion models to 8B parameters and achieves results similar to LLaMA3-8B across both pretraining and instruction-tuned evaluations. Furthermore, Nie et al. (nie2024scaling, ) provide scaling law analysis showing that diffusion models follow similar power-law trends as AR models, though they may require up to 16$\times$ more compute under single-epoch training, Swerdlow et al. (swerdlow2025unidisc, ) find similar trends on multimodal data containing both image and text. However, these evaluations are restricted to single-pass training and do not examine the data-constrained, multi-epoch regimes which is the focus of our work.
3 Method
Our objective is to determine whether masked diffusion language models are more effective than standard autoregressive models in data-constrainted settings. The main difference between AR and diffusion models is the way they factorize the joint distribution of the sequence. Masked diffusion factorizes the joint distribution of the sequence in a random order, while AR factorizes the joint distribution of the sequence in a left-to-right order. To isolate the impact of this, we keep the architecture and data pipeline fixed across both families and vary only the factorization of the joint distribution.
3.1 Preliminaries:
Autoregressive models.
In Autoregressive LLMs (touvron2023llama, ; radford2019language, ; brown2020language, ) we predict each token based on a growing prefix of prior tokens, defining a left-to-right factorization of the sequence probability:
$p_{\text{AR}}(x_{1},\dots,x_{L})=\prod_{j=1}^{L}p(x_{j}\mid x_{<j}).$
This structure is implemented using a causal attention mask, which prevents each token from attending to future positions. The model is trained via next-token prediction over clean, uncorrupted sequences.
Diffusion models.
Masked diffusion language models  (austin2021structured, ; sahoo2024simple, ; nie2024scaling, ; swerdlow2025unidisc, ) treat generation as iterative denoising. For each training sequence $x=(x_{1},\dots,x_{L})$ we
1. Corrupt the sequence by sampling a masking ratio $r\sim\mathcal{U}(0,1)$ and independently replacing each token with a special [MASK] symbol with probability $r$. This yields a corrupted sequence $\tilde{x}$ and a mask set
[…]
2. Denoise by predicting the original tokens at the masked positions with full (bidirectional) attention over $\tilde{x}$:
$p_{\text{Diffusion}}(x\mid\tilde{x})\;=\;\prod_{i\in\mathcal{M}}p_{\theta}\!\bigl{(}x_{i}\,\mid\,\tilde{x}\bigr{)}.$
Because the mask pattern is resampled for every example, the model is implicitly trained on a vast collection of token–ordering tasks; the standard left-to-right ordering used by AR models is just one ordering within this ensemble. The absence of a causal mask allows each prediction to attend to both past and future unmasked tokens, making the factorization fundamentally non-sequential.
3.2 Modeling Details for AR and Masked Diffusion
Our goal is to isolate the impact of the factorization—fixed left-to-right versus random-order denoising—while keeping every other design choice constant. Unless noted otherwise, both model families share the same Transformer backbone (GPT-2 style with rotary positional embeddings, RoPE (su2021roformer, )) and identical training hyper-parameters across the full parameter sweep (7 M – 2 B).
Given a clean input sequence $x=(x_{1},\dots,x_{L})\in\mathcal{V}^{L}$, both models minimize a token-level cross-entropy loss, yet they differ in the conditioning context:
Autoregressive (AR) objective.
AR models predict each token conditioned on its prefix using a causal attention mask:
$\mathcal{L}_{\text{AR}}\;=\;-\sum_{j=2}^{L}\log p_{\theta}\!\bigl{(}x_{j}\mid x_{<j}\bigr{)}.$
Masked Diffusion objective.
[…]
$\mathcal{L}_{\text{Diffusion}}\;=\;-\mathbb{E}_{r}\;\mathbb{E}_{\tilde{x}\sim q_{r}}\;\sum_{i\in\mathcal{M}}\log p_{\theta}\!\bigl{(}x_{i}\mid\tilde{x}\bigr{)}.$
Unlike Nie et al. (nie2024scaling, ), we do not add auxiliary losses (e.g., classifier-free guidance) in order to ablate only the factorization.
Beyond the attention mechanism and input corruption, all other variables are held constant. We follow the hyperparameter configuration proposed by Muennighoff et al. muennighoff2023scaling for all training runs. In particular, we use a dynamic learning rate schedule that adapts to the number of training epochs. The only distinctions between AR and diffusion models in our implementation are:
  1. 1.
    Attention mechanism: Causal attention for AR; full self-attention for masked diffusion.
  2. 2.
    Prediction target: AR models predict the next token; diffusion models predict the masked tokens.
3.3 Scaling Framework in Data-Constrained Settings
Classical scaling laws, such as those proposed by (kaplan2020scaling, ; hoffmann2022chinchilla, ), model validation loss as a function of total parameters ($N$) and training tokens ($D$), assuming all data is unique. These laws have been instrumental in guiding compute-optimal training of language models. However, this assumption becomes unrealistic as the community approaches the limits of high-quality text data available on the internet.
To address this, Muennighoff et al.muennighoff2023scaling extend the Chinchilla framework to explicitly account for repeated data — a common necessity in data-constrained regimes. They show that repeating training data beyond a few epochs yields diminishing returns and propose a new scaling law that incorporates the decaying utility of repeated tokens.
We briefly outline their formulation below.
Definitions:
  • $U$: number of unique tokens available for training,
  • $E$: number of epochs (i.e., how many times each unique token is reused),
  • $D=U\cdot E$: total number of tokens seen by the model.
To model diminishing returns from repeated data, Muennighoff et al. muennighoff2023scaling introduce an effective unique data size $D^{\prime}$, motivated by the idea that each additional epoch contributes less useful signal than the previous. Specifically, they assume the value extracted from the $k^{\text{th}}$ exposure to the same data follows a geometric progression, where the utility of a token on its $k$-th repetition is $(1-\delta)^{k-1}$. Summing over all epochs the total effective data becomes: $D^{\prime}=U\cdot\sum_{k=1}^{E}(1-\delta)^{k-1}=U\cdot\frac{1-(1-\delta)^{E}}{\delta}$ where $\delta$ is the decay factor. Defining $R_{D}^{\star}=\frac{1-\delta}{\delta}$, the expression simplifies to the exponential-decay form:
$D^{\prime}=U+U\cdot R_{D}^{\star}\left(1-e^{-(E-1)/R_{D}^{\star}}\right).$
here $R_{D}^{*}$ represents the half-life of data reuse, repeating data beyond $R_{D}^{*}$ epochs will result in significant diminishing returns. This form approximates the geometric sum well and captures diminishing returns over repeated epochs. As the number of epochs $E\to\infty$, the exponential term vanishes and $D^{\prime}$ asymptotically approaches: $D^{\prime}\to U+U\cdot R_{D}^{\star},$ implying that no matter how many times data is repeated, the maximum usable signal is bounded by $(1+R_{D}^{\star})\cdot U$. This defines a natural saturation point on returns: even infinite compute yields no additional effective data beyond this limit.
A symmetric formulation is applied to model parameters for mathetmatical convenience which is used to define $N^{\prime}$. Finally, a modified Chinchilla-style loss function incorporates these effective quantities $N^{\prime}$ and $D^{\prime}$:
$\mathcal{L}(N,D)=\frac{A}{(N^{\prime})^{\alpha}}+\frac{B}{(D^{\prime})^{\beta}}+E_{0},$
with $A,B,\alpha,\beta,E_{0},R_{D}^{\star},N_{D}^{\star}$ fitted empirically from training runs. This formulation accurately captures loss behavior in regimes where data is reused multiple times and serves as a powerful tool for guiding training under data scarcity.
In this work, we adopt this framework to study how diffusion models and autoregressive models compare in their ability to extract value from repeated data, enabling apples-to-apples comparisons across compute, data, and model scale.
3.4 Training setup
We use the English C4 corpus (raffel2020exploring, ), tokenized with the GPT-2 BPE vocabulary and truncated or padded to 2048 tokens per sequence. We consider unique-token budgets of $U\in\{25,50,100\}$M and train for up to 800 epochs (80B tokens total). Models are trained ranging from 7M to 2.5B parameters, following the Chinchilla scaling strategy where both width and depth are increased proportionally. The detailed architectural configurations of each model are provided in Appendix 10. For all training runs, we adopt the hyperparameter configuration introduced by Muennighoff et al. muennighoff2023scaling . This may provide a slight advantage to autoregressive models, as these hyperparameters were originally tuned for that family. For all models, we use the following hyperparameters: batch size of 256 sequences, AdamW optimizer with $\beta_{1}{=}0.9$, $\beta_{2}{=}0.95$, $\epsilon{=}10^{-8}$, a learning rate schedule with peak 2e-4, minimum 2e-5, 1% warm-up, cosine decay, weight decay 0.1, and gradient clipping of 1.0.
4 Experiments
Our goal is to compare the performance of masked diffusion models and autoregressive models in data-constrained settings. To this end, we train a total of 200 models—100 diffusion models and 100 autoregressive models—across varying unique data sizes, model scales, and epoch counts. We present the empirical results in Section 4.1. In Section 4.2, we fit scaling laws tailored to data-constrained regimes for both model types, following the methodology introduced by Muennighoff et al.muennighoff2023scaling . These scaling laws allow us to analyze performance trends and identify scenarios where diffusion models should be preferred over autoregressive ones (Section 4.3). Finally, in Section 4.4, we demonstrate that the superior validation loss of diffusion models indeed correlates with improved downstream task performance.
4.1 Does Diffusion Beat AR in Data-Constrained Settings?
Prior comparisons between diffusion and autoregressive (AR) language models have largely focused on the single-epoch regime, where each token is seen only once during training mdlm ; nie2024scaling ; swerdlow2025unidisc . In this setting, diffusion models are consistently reported to require substantially more training compute ($C\sim 6ND$) than AR models to achieve comparable validation loss. For instance, Nie et al. (nie2024scaling, ) and Swerdlow et al. (swerdlow2025unidisc, ) derive scaling laws showing that masked diffusion models can require up to 16$\times$ more compute than AR counterparts.
Crucially, these studies scale compute by increasing both the model size ($N$) and the amount of unique training data ($D$) proportionally. As a result, they do not isolate whether diffusion’s 16x inefficiency stems from needing more total compute—or more unique data.
In other words: is diffusion limited by compute efficiency or by data efficiency?
To answer this, we systematically study diffusion models in data-constrained settings, where the total amount of unique data is fixed and models are trained for many epochs, reusing the same data. Unlike prior work, our evaluation explicitly decouples model scaling from data reuse, allowing us to disentangle the effects of compute and data.
We train a large suite of AR and diffusion models across three unique data regimes—25M, 50M, and 100M tokens—and a wide range of training compute budgets. In Figure 1, we report empirical validation loss as a function of training FLOPs for the 50M and 100M regimes; results for the 25M setting are shown in Appendix Figure 7. We find that AR models initially outperform diffusion models when trained with the compute-optimal budget prescribed by Chinchilla scaling laws (denoted by the solid vertical line). However, this advantage disappears as training continues beyond this point. When models are allowed to train for additional epochs on repeated data, diffusion models consistently surpass AR models in validation loss across all data regimes. These findings indicate that the previously observed inefficiency of diffusion models is largely a consequence of evaluating them solely in the single-epoch regime. In data-constrained settings with repeated exposures, diffusion models extract significantly more value from the same data than their AR counterparts.
A key question remains is how should one go about increasing compute for diffusion models: by increasing model size, or by increasing the number of epochs (i.e., data reuse)? To address this, we analyze the trade-off between parameters and epochs in Figure 2, which shows validation loss contours as a function of both axes. In the 100M unique token regime, for example, we find that diffusion achieves its best loss at 500 epochs, while AR model reach its best at just 50 epochs. Each point on the contour plot corresponds to a model trained with a specific parameter count and number of epochs; we report the actual validation loss at each configuration, without early stopping. We find that autoregressive models begin to overfit at high epoch counts, with validation loss worsening as training continues beyond a certain point. In contrast, diffusion models show no signs of overfitting within our compute budget—the best validation loss is achieved at the highest epoch counts we explore. This suggests that diffusion models continue to benefit from additional training on repeated data, and that observing overfitting may require significantly more compute.
To contextualize these results, we highlight two key configurations in Figure 2 for each model family: the compute-optimal point for single-epoch training, as identified by prior scaling law analyses hoffmann2022chinchilla ; nie2019adversarial (marked with a colored star in the bottom-left), and the best validation loss achieved under extended multi-epoch training (marked with a black star). At the compute-optimal point, which corresponds to training for a single epoch, diffusion models perform substantially worse than autoregressive models (10.65 vs. 7.07), consistent with prior findings that diffusion performs worse initially. However, as training is extended to hundreds of epochs, diffusion models continue to improve and eventually achieve a lower validation loss (3.55) than the best AR models (3.71). While AR models begin to overfit as training progresses, diffusion models show no signs of overfitting within our budget.
(a) Autoregressive contour: validation loss over epochs and model sizes.
(b) Diffusion contour: validation loss over epochs and model sizes.
Figure 2: Validation loss contours over epochs and model sizes for autoregressive (left) and diffusion (right) models, trained on 100M unique tokens. Each plot shows validation loss as a function of training epochs (x-axis) and model parameters (y-axis). The colored star marks the compute-optimal point for single-epoch training, as predicted by prior scaling laws hoffmann2022chinchilla ; nie2019adversarial , and the black star indicates the lowest validation loss achieved through extended multi-epoch training. In the single-epoch regime, diffusion models perform worse than AR models (10.65 vs. 7.07). However, when trained longer, diffusion models achieve a substantially lower final loss (3.55 vs. 3.71). This corresponds to a 67% reduction in loss for diffusion models compared to just 48% for AR models, highlighting their superior ability to leverage repeated data. These results underscore that diffusion models require significantly more training—both in epochs and compute—to realize their advantages in data-constrained settings.
4.2 Fitting Data-Constrained Scaling Laws
To gain deeper insight into the trade-offs between diffusion and autoregressive models in data-constrained settings, we fit scaling laws to both model families across single-epoch and multi-epoch regimes, as described in Section 3.3. Our approach systematically varies three key factors: (1) the amount of unique data, (2) model parameter count, and (3) number of training epochs. This grid search allows us to disentangle the effects of data quantity, model capacity, and data reuse on final model performance.
We evaluate the quality of our scaling law fits using the coefficient of determination ($R^{2}$) and relative prediction error, as shown in Table 1. For autoregressive models, our $R^{2}$ values closely match those reported by Muennighoff et al. muennighoff2023scaling , indicating consistent behavior under repeated training. Interestingly, diffusion models yield significantly higher $R^{2}$ values, reflecting a better overall fit. We attribute this to lower variance in validation loss across training runs, likely due to the absence of overfitting in diffusion models even at high epoch counts.
Beyond the overall fit, we extract two key parameters from the scaling laws: $R_{D}^{*}$, which characterizes the effective half-life of data reuse—i.e., the number of epochs after which additional training on repeated data yields diminishing returns—and $R_{N}^{*}$, which indicates the optimal model size for a given data budget. Our results reveal a sharp contrast in data reuse half-lives: diffusion models exhibit an $R_{D}^{*}$ of 512.85, compared to just 31.93 for autoregressive models. A higher $R_{D}^{*}$ implies that a model can benefit from many more repeated exposures before saturating. This suggests that diffusion models continue to improve across hundreds of epochs, while AR models quickly saturate—highlighting the superior data efficiency of diffusion models in data-constrained regimes.
Table 1: Fitting metrics of the scaling law model for Diffusion and AR. Diffusion and AR achieve a strong fit across both phases.
(a) Initial fit.
Model $R^{2}$ Loss
Diffusion 0.9447 $0.0002$
AR 0.9439 $7.7532\text{e}{-05}$
(b) Second step fit with extracted scaling parameters.
Model $R^{2}$ Loss $R_{D}^{*}$ $R_{N}^{*}$
Diffusion 0.9784 0.00079 493.89 1265.65
AR 0.7628 0.00361 31.19 55.16
Figure 3 illustrates how the utility of unique data decays with increased repetition. We evaluate this effect across three compute budgets—$1\times 10^{19}$, $3\times 10^{19}$, and $1\times 10^{20}$ FLOPs—by varying the proportion of unique data and parameters while keeping total compute fixed (e.g., 50% of the data for 2 epochs, 25% for 4 epochs, etc.). For each compute budget, we use single-epoch scaling laws to determine the optimal model size and unique token count for both AR and diffusion models. This experimental design allows us to directly measure how the utility of data diminishes with increased repetition. We present both empirical results and fitted curves from our parametric scaling law, observing strong agreement between the two. Notably, the decay rate of data value remains consistent across compute budgets for both model families. However, diffusion models consistently exhibit a substantially slower decay rate than AR models, suggesting they are better able to extract value from repeated data.
Figure 4 shows validation loss versus training tokens and reinforces the trend: AR models overfit with increased repetition, showing diverging loss curves. In contrast, diffusion models exhibit overlapping curves across repetitions, indicating no signs of overfitting and a very low decay rate with data reuse..
Figure 5 shows extrapolated training curves at large compute budgets for AR and diffusion models using our fitted scaling laws. For each setting, we use the compute-optimal model and dataset size derived from single-epoch scaling laws and extend training across multiple epochs. The dashed lines represent the ideal Chinchilla-style scaling behavior, where all training tokens are assumed to be unique. We find that for AR models, repeated data provides nearly the same benefit as fresh data only up to about 4 epochs. Beyond this point, additional repetition yields diminishing returns. In contrast, diffusion models continue to match the unique-data curve for up to 100 epochs, indicating a far greater capacity to benefit from repeated data in data-constrained regimes.
Figure 3: Decay rate of data value under repetition: left shows diffusion, middle AR, and right the average decay rate for both. Points are empirical results (darker color = higher FLOPs, lighter color = lower FLOPs; each line = fixed compute), we find that fitted curves (represented as lines) closely match the empirical points, indicating our scaling laws are representative. The decay rate of value for repeated data is lower for diffusion, reflecting its greater robustness to repeating.
(a) AR Training Curves.
(b) Diffusion Training Curves.
Figure 4: Training curves for different epoch counts, all with using the same total compute. Each curve shows a different tradeoff between unique data and repetition. For AR models, validation loss rises with more epochs (overfitting), while for diffusion models, the curves are nearly unchanged, showing much greater robustness to data repetition.
Figure 5: Predicted validation loss for AR models (left) and Diffusion models (right) under compute-optimal settings, extrapolated to larger compute budgets. Dotted lines indicate the hypothetical case where repeated data is as valuable as new data. For AR, this holds up to about 4 epochs; for diffusion, up to 100 epochs, showing that diffusion models are much more robust to data repetition.
4.3 When to Use Diffusion over AR?
A key question for practitioners is: when should diffusion be preferred over autoregressive models (AR)? To answer this, we compare the fitted data-constrained scaling laws for both model families (§3.3).
We define the validation loss gap between diffusion and AR as:
$\Delta\mathcal{L}(C,U)=\mathcal{L}_{\text{Diffusion}}(C,U)-\mathcal{L}_{\text{AR}}(C,U),$
where $C$ is total training compute and $U$ is the number of unique tokens. Positive values favor AR; negative values favor diffusion. The critical compute $C_{\mathrm{crit}}(U)$ is the point where the models perform equally:
$\Delta\mathcal{L}(C_{\mathrm{crit}},U)=0.$
Figure 6(a) shows a heatmap of $\Delta\mathcal{L}$ over compute and data. Red regions indicate regimes where diffusion outperforms AR ($\Delta\mathcal{L}<0$), while blue regions favor AR. As expected, AR performs better in low-compute settings due to its efficient per-step learning. However, diffusion models begin to outperform AR at higher compute, especially when data is limited and repeated.
Figure 6(b) plots the critical compute frontier $C_{\mathrm{crit}}(U)$—the compute required for diffusion to match AR at a given unique token count $U$. This frontier follows a power law:
$C_{\mathrm{crit}}(U)\propto U^{2.174}.$
The linear fit in log-log space is:
$\log_{10}(U)=0.460\cdot\log_{10}(C)-7.052,\quad\text{so}\quad C_{\text{crit}}(U)=2.12\times 10^{15}\cdot U^{2.174}.$
The green dashed line shows the fitted curve, and the blue crosses represent empirical crossover points—where diffusion matches AR performance in experiments. These points align closely with the predicted frontier, confirming our fitted equation’s accuracy.
(a) Loss Gap Heatmap. Difference in validation loss ($\Delta\mathcal{L}=\mathcal{L}_{\text{Diffusion}}-\mathcal{L}_{\text{AR}}$) across unique data sizes and FLOPs. Red indicates regions where diffusion outperforms AR models and blue where AR outperforms diffusion.
(b) Critical Compute Curve. The FLOPs threshold $C_{\mathrm{crit}}(U)$ beyond which diffusion outperforms AR models. This follows a power law: $C_{\mathrm{crit}}(U)\propto U^{2.174}$.
Figure 6: When does Diffusion beat AR? Left: Heatmap showing where diffusion models have lower validation loss than AR models. Right: The critical compute curve defining the compute threshold needed for diffusion to match autoregressive models at a given unique token count.
4.4 Downstream Results
We evaluate the best-performing diffusion and autoregressive (AR) models on several downstream benchmarks to assess whether the gains in validation loss translate to practical improvements in generalization.
Motivated by the critical compute threshold equation identified in Section 4.3, we scale the training data to 500M unique tokens and train a 2.3B parameter diffusion model using the compute budget predicted by the critical compute limit. The model was trained for 130 epochs, during which we observed no signs of convergence. Training was terminated due to compute constraints.
Across a diverse set of tasks and data scales, diffusion models consistently achieve higher accuracy than their AR counterparts. This validates our findings in Section 4.3, also confirms that the data efficiency gains observed in validation loss translate into stronger downstream performance. Additional results, including performance on textbook-style datasets, are provided in Table 3 in the Appendix.
Table 2: Downstream Results for the best autoregressive and diffusion trained in different data-constrainted settings. We report the results for the models with the best validation loss in 100M and 500M unique data regime.To better understand the difficulty of each benchmark we also report the accuracy of random baseline
Benchmarks Random Baseline 100M unique tokens 500M unique tokens
AR Diffusion AR Diffusion
ARC-Easy clark2018think 25.00 35.63 37.84 43.79 45.95
BoolQ clark2019boolq 50.00 46.00 49.38 51.87 55.26
COPA roemmele2011choice 50.00 56.33 59.00 67.00 64.83
HellaSwag zellers2019hellaswag 25.00 27.37 30.24 32.28 35.33
PiQA 50.00 60.94 60.72 65.71 65.61
RACE lai2017race 25.00 25.28 28.96 28.28 31.44
WinoGrande XL sakaguchi2021winogrande 50.00 48.87 50.97 50.61 51.51
SciQ SciQ 25.00 58.05 68.67 67.82 79.13
Lambada paperno2016lambada 00.00 10.91 15.19 15.07 22.30
Note: All values represent accuracy (%). Best results shown in bold.
5 Discussion
Why do diffusion models outperform autoregressive (AR) models in low-data regimes?
We hypothesize that the key advantage stems from the use of random masking in diffusion models, which serves as a form of data augmentation. Unlike AR models, which are trained on a single, fixed left-to-right ordering, masked diffusion models are exposed to a wide variety of token prediction tasks and orderings during training—including, the left-to-right order itself. This broader distribution over prediction tasks encourages better generalization and more effective use of each training example. Analogous to image-based tasks where techniques like random cropping or color jittering boost generalization  caron2021emerging ; cubuk2020randaugment ; he2022masked , we believe the corruption process in diffusion plays a similar role. By learning to denoise across many corruptions, the model extracts richer signal per example over time, resulting in improved data efficiency.
Why are autoregressive (AR) models more compute-efficient than diffusion models?
We hypothesize two main contributing factors. (i) Order specialization: AR models are trained with a fixed left-to-right factorization, so every gradient update reinforces the same prediction task, allowing them to specialize effectively. In contrast, diffusion models must generalize across many random token orderings, which hinders specialization. (ii) Stronger supervision per update: In AR training, every token in a training sequence serves as a supervised target, and the causal structure enables dense gradient updates, resulting in stable, low-variance learning. Diffusion models, however, compute loss only on a subset of masked tokens, making supervision sparser per sequence, even though gradients propagate through the entire input. As a result, each update carries less direct learning signal. Arriola et al. (arriola2025block, ) show that tuning the masking schedule can help reduce gradient variance and improve training compute efficiency.
6 Limitations
In this work, we examined two extremes of generative modeling: masked diffusion models, which learn random token orderings and are more data-efficient, and autoregressive (AR) models, which follow a fixed left-to-right order and are more compute-efficient. While our results highlight a clear trade-off, this need not be binary—hybrid models that interpolate between AR and diffusion would offer a better balance. Although prior works have explored such hybrids (arriola2025block, ; hoogeboom2021autoregressive, ), they have not been evaluated through the lens of data-compute efficiency. Additionally, our scaling laws are currently fit over a limited range of unique data sizes; extending them to larger regimes may improve predictive accuracy and reveal further insights.
7 Conclusion
As the availability of high-quality data plateaus, improving data efficiency becomes essential for scaling deep learning. In this work, we show that masked diffusion models consistently outperform autoregressive (AR) models in data-constrained regimes — when training involves repeated passes over a limited dataset. We establish new scaling laws for diffusion models, revealing their ability to extract value from repeated data far beyond what AR models can achieve. These results challenge the conventional belief that AR models are universally superior and highlight diffusion models as a compelling alternative when data—not compute—is the primary bottleneck. Looking ahead, efficient use of finite data may define the next frontier in scaling deep learning models. Although the studies have been performed in the context of language models, we believe these findings should apply across any kind of sequence modeling data, such as in robotics or healthcare.
For practitioners, our takeaway is simple: if you are compute-constrained, use autoregressive models; if you are data-constrained, use diffusion models.
8 Acknowledgement
We thank Alexander Li for his valuable insights and detailed feedback on the manuscript. We are also grateful to Zheyang Qin for his help in improving the figures. Finally, we thank Niklas Muennighoff, Simo Ryu and Colin Raffel for their helpful comments on the final draft of the paper. This work was supported in part by ONR MURI N00014-22-1-2773, ONR N00014-22-1-2096, and AFOSR FA9550-23-1-0747.

### v2

@@
 However, diffusion models consistently exhibit a substantially slower decay rate than AR models, suggesting they are better able to extract value from repeated data.
-Figure 4 shows validation loss versus training tokens and reinforces the trend: AR models overfit with increased repetition, showing diverging loss curves.
-In contrast, diffusion models exhibit overlapping curves across repetitions, indicating no signs of overfitting and a very low decay rate with data reuse..
-Figure 5 shows extrapolated training curves at large compute budgets for AR and diffusion models using our fitted scaling laws.
-For each setting, we use the compute-optimal model and dataset size derived from single-epoch scaling laws and extend training across multiple epochs.
+Figure 4 shows validation loss versus training tokens using the compute budget of 1e19.
+The results reinforces the trend: AR models overfit with increased repetition, showing diverging loss curves.
+In contrast, diffusion models exhibit overlapping curves across repetitions, indicating no signs of overfitting and a very low decay rate with data reuse.
+Figure 5 shows extrapolated training curves at large compute budgets.
+For each setting, we use the compute-optimal model and dataset size derived from single-epoch scaling laws for 1e19, 3e19 and 1e20.
+We then extend training to multiple epochs.
 The dashed lines represent the ideal Chinchilla-style scaling behavior, where all training tokens are assumed to be unique.
@@
 The linear fit in log-log space is:
-$\log_{10}(U)=0.460\cdot\log_{10}(C)-7.052,\quad\text{so}\quad C_{\text{crit}}(U)=2.12\times 10^{15}\cdot U^{2.174}.$
+$\log_{10}(U)=0.460\cdot\log_{10}(C)-7.052,\quad\text{so}\quad C_{\text{crit}}(U)=2.12\times 10^{1.956}\cdot U^{2.174}.$
 The green dashed line shows the fitted curve, and the blue crosses represent empirical crossover points—where diffusion matches AR performance in experiments.
@@
 Why do diffusion models outperform autoregressive (AR) models in low-data regimes?
-We hypothesize that the key advantage stems from the use of random masking in diffusion models, which serves as a form of data augmentation.
+We hypothesize that the key advantage stems from the use of random factorization in diffusion models, which serves as a form of data augmentation.
 Unlike AR models, which are trained on a single, fixed left-to-right ordering, masked diffusion models are exposed to a wide variety of token prediction tasks and orderings during training—including, the left-to-right order itself.
@@
 Finally, we thank Niklas Muennighoff, Simo Ryu and Colin Raffel for their helpful comments on the final draft of the paper.
+We thank Stella Biderman for pointing out the inconsistent unit used in the Critical Compute equation.
 This work was supported in part by ONR MURI N00014-22-1-2773, ONR N00014-22-1-2096, and AFOSR FA9550-23-1-0747.

### v3

@@
 […]
-$\mathcal{L}_{\text{Diffusion}}\;=\;-\mathbb{E}_{r}\;\mathbb{E}_{\tilde{x}\sim q_{r}}\;\sum_{i\in\mathcal{M}}\log p_{\theta}\!\bigl{(}x_{i}\mid\tilde{x}\bigr{)}.$
-Unlike Nie et al. (nie2024scaling, ), we do not add auxiliary losses (e.g., classifier-free guidance) in order to ablate only the factorization.
+$\mathcal{L}_{\text{Diffusion}}\;=\;-\mathbb{E}_{r}\;\mathbb{E}_{\tilde{x}\sim q_{r}}\;\frac{1}{r}\sum_{i\in\mathcal{M}}\log p_{\theta}\!\bigl{(}x_{i}\mid\tilde{x}\bigr{)}.$
 Beyond the attention mechanism and input corruption, all other variables are held constant.
@@
 Model $R^{2}$ Loss
-Diffusion 0.9447 $0.0002$
-AR 0.9439 $7.7532\text{e}{-05}$
+Diffusion 0.9447 0.00020.00020.0002
+[…]
 (b) Second step fit with extracted scaling parameters.
@@
 AR 0.7628 0.00361 31.19 55.16
-Figure 3 illustrates how the utility of unique data decays with increased repetition.
-We evaluate this effect across three compute budgets—$1\times 10^{19}$, $3\times 10^{19}$, and $1\times 10^{20}$ FLOPs—by varying the proportion of unique data and parameters while keeping total compute fixed (e.g., 50% of the data for 2 epochs, 25% for 4 epochs, etc.).
-For each compute budget, we use single-epoch scaling laws to determine the optimal model size and unique token count for both AR and diffusion models.
-This experimental design allows us to directly measure how the utility of data diminishes with increased repetition.
-We present both empirical results and fitted curves from our parametric scaling law, observing strong agreement between the two.
-Notably, the decay rate of data value remains consistent across compute budgets for both model families.
-However, diffusion models consistently exhibit a substantially slower decay rate than AR models, suggesting they are better able to extract value from repeated data.
+[…]
 Figure 4 shows validation loss versus training tokens using the compute budget of 1e19.
@@
 The linear fit in log-log space is:
-$\log_{10}(U)=0.460\cdot\log_{10}(C)-7.052,\quad\text{so}\quad C_{\text{crit}}(U)=2.12\times 10^{1.956}\cdot U^{2.174}.$
+$\log_{10}(U)=0.460\cdot\log_{10}(C)-1.050,\quad\text{so}\quad C_{\text{crit}}(U)=2.12\times 10^{1.956}\cdot U^{2.174}.$
 The green dashed line shows the fitted curve, and the blue crosses represent empirical crossover points—where diffusion matches AR performance in experiments.
@@
 Motivated by the critical compute threshold equation identified in Section 4.3, we scale the training data to 500M unique tokens and train a 2.3B parameter diffusion model using the compute budget predicted by the critical compute limit.
-The model was trained for 130 epochs, during which we observed no signs of convergence.
-Training was terminated due to compute constraints.
-Across a diverse set of tasks and data scales, diffusion models consistently achieve higher accuracy than their AR counterparts.
+We train the model for 130 epochs, during which we observe no signs of convergence.
+We terminate the training due to compute constraints.
+Across a diverse set of tasks and data scales, diffusion models consistently outperform their AR counterparts.
 This validates our findings in Section 4.3, also confirms that the data efficiency gains observed in validation loss translate into stronger downstream performance.
@@
 Why do diffusion models outperform autoregressive (AR) models in low-data regimes?
-We hypothesize that the key advantage stems from the use of random factorization in diffusion models, which serves as a form of data augmentation.
+We hypothesize that the key advantage stems from the random factorizations in diffusion objective, which serves as a form of implicit data augmentation.
 Unlike AR models, which are trained on a single, fixed left-to-right ordering, masked diffusion models are exposed to a wide variety of token prediction tasks and orderings during training—including, the left-to-right order itself.
@@
 By learning to denoise across many corruptions, the model extracts richer signal per example over time, resulting in improved data efficiency.
+We further analyze this question in Section 9 in Appendix.
 Why are autoregressive (AR) models more compute-efficient than diffusion models?

### v4

@@
 Furthermore, there are several domains, such as robotics and healthcare, where the data, not compute, is a scarce resource even to begin with.
-LLM development has so far been dominated by autoregressive (AR) models, which factorize the joint distribution of text in a fixed left-to-right order, i.e., $p(x)=p(x_{0})p(x_{1}\mid x_{0})p(x_{2}\mid x_{1})p(x_{3}\mid x_{1},x_{2})$.
+LLM development has so far been dominated by autoregressive (AR) models, which factorize the joint distribution of text in a fixed left-to-right order.
 While this modeling approach has delivered state-of-the-art performance across a range of benchmarks, it remains unclear whether it is the optimal strategy going forward.
-Recently, diffusion-based models—specifically masked diffusion models  austin2021structured ; sahoo2024simple ; lou2310discrete ; shi2024simplified ; arriola2025block —have emerged as an alternative strategy, where they model text via a random-order factorization, e.g., $p(x)=p(x_{2}|x_{3})p(x_{3}|x_{0})p(x_{0}|x_{1})p(x_{1})$.
-These models learn across many different masked versions of the text, allowing them to train on diverse token orderings.
-Although diffusion language models have demonstrated similar scaling behavior to AR models (nie2024scaling, ; swerdlow2025unidisc, ), their practical benefits have, so far, been modest—largely due to their high training compute requirements.
+Recently, diffusion-based models—specifically masked diffusion models  austin2021structured ; sahoo2024simple ; lou2310discrete ; shi2024simplified ; arriola2025block —have emerged as an alternative strategy, framing text generation as an iterative masked denoising process rather than next-token prediction.
+At each step, the model predicts a randomly masked subset of tokens conditioned on the remaining ones, implicitly averaging over many conditional prediction orders instead of committing to one.
+Although these models have demonstrated similar scaling behavior to AR models (nie2024scaling, ; swerdlow2025unidisc, ), their practical benefits have, so far, been modest—largely due to their high training compute requirements.
 This high compute demand has become the central obstacle to wider adoption of diffusion-based language models.
@@
 4.1 Does Diffusion Beat AR in Data-Constrained Settings?
-Prior comparisons between diffusion and autoregressive (AR) language models have largely focused on the single-epoch regime, where each token is seen only once during training mdlm ; nie2024scaling ; swerdlow2025unidisc .
+Prior comparisons between diffusion and autoregressive (AR) language models have largely focused on the single-epoch regime, where each token is seen only once during training nie2024scaling ; swerdlow2025unidisc .
 In this setting, diffusion models are consistently reported to require substantially more training compute ($C\sim 6ND$) than AR models to achieve comparable validation loss.

### v5

@@
 muennighoff2023scaling fit scaling laws for AR models in data-constrainted settings and define $R_{D}^{*}$ as a learned constant that characterizes the number of epochs after which training more epochs results in significantly diminished returns.
-For autoregressive models, they estimate $R_{D}^{*}\approx 15$ In contrast, we find $R_{D}^{*}\approx 500$ for diffusion models, suggesting they can benefit from repeated data over far more epochs without major degradation (Section 4.2).
+For autoregressive models, they estimate $R_{D}^{*}\approx 15$ .
+In contrast, we find $R_{D}^{*}\approx 500$ for diffusion models, suggesting they can benefit from repeated data over far more epochs without major degradation (Section 4.2).
   4. 4.
@@
 Diffusion-Based Language Models.
-Diffusion models, originally developed for image generation (ho2020denoising, ), have recently been adapted to text, offering a fundamentally different paradigm for language modeling (austin2021structured, ; li2022diffusion, ; gulrajani2023likelihood, ).
-Broadly, diffusion language models fall into two categories: continuous and discrete.
-Continuous approaches (gulrajani2023likelihood, ) inject Gaussian noise in the forward process, whereas discrete methods (austin2021structured, ) corrupt tokens with noise sampled from distributions such as Bernoulli.
-Among the two classes, continuous diffusion has proven more difficult to scale on language data gulrajani2023likelihood ; lou2023discrete .
-In contrast, recent advances in discrete diffusion—particularly masked diffusion—have shown encouraging results.
-Recent work (arriola2025block, ; dieleman2023language, ; sahoo2024simple, ; lou2023discrete, ) has significantly narrowed the performance gap between diffusion and AR models.
-Notably, LLaDA (nie2025large, ) scales masked diffusion models to 8B parameters and achieves results similar to LLaMA3-8B across both pretraining and instruction-tuned evaluations.
-Furthermore, Nie et al.
-(nie2024scaling, ) provide scaling law analysis showing that diffusion models follow similar power-law trends as AR models, though they may require up to 16$\times$ more compute under single-epoch training, Swerdlow et al.
-(swerdlow2025unidisc, ) find similar trends on multimodal data containing both image and text.
-However, these evaluations are restricted to single-pass training and do not examine the data-constrained, multi-epoch regimes which is the focus of our work.
+[…]
 3 Method
@@
 This yields a corrupted sequence $\tilde{x}$ and a mask set
-[…]
+$\mathcal{M}\;=\;\{\,i\in[1,L]:\tilde{x}_{i}=\textsc{[MASK]}\,\}.$
 2. Denoise by predicting the original tokens at the masked positions with full (bidirectional) attention over $\tilde{x}$:
@@
 Masked Diffusion objective.
-[…]
+For masked diffusion we first sample a masking ratio $r\sim\mathcal{U}(0,1)$ and construct a corrupted sequence $\tilde{x}$ by independently replacing each token with [MASK] with probability $r$.
+Let $\mathcal{M}=\{\,i:\tilde{x}_{i}=\textsc{[MASK]}\,\}$ be the set of masked positions.
+The loss is then
 $\mathcal{L}_{\text{Diffusion}}\;=\;-\mathbb{E}_{r}\;\mathbb{E}_{\tilde{x}\sim q_{r}}\;\frac{1}{r}\sum_{i\in\mathcal{M}}\log p_{\theta}\!\bigl{(}x_{i}\mid\tilde{x}\bigr{)}$.
@@
 4.1 Does Diffusion Beat AR in Data-Constrained Settings?
-Prior comparisons between diffusion and autoregressive (AR) language models have largely focused on the single-epoch regime, where each token is seen only once during training nie2024scaling ; swerdlow2025unidisc .
-In this setting, diffusion models are consistently reported to require substantially more training compute ($C\sim 6ND$) than AR models to achieve comparable validation loss.
-For instance, Nie et al.
-(nie2024scaling, ) and Swerdlow et al.
-(swerdlow2025unidisc, ) derive scaling laws showing that masked diffusion models can require up to 16$\times$ more compute than AR counterparts.
+[…]
 Crucially, these studies scale compute by increasing both the model size ($N$) and the amount of unique training data ($D$) proportionally.
@@
 Model $R^{2}$ Loss
-Diffusion 0.9447 0.00020.00020.0002
-[…]
+Diffusion 0.9447 $0.0002$
+AR 0.9439 $7.7532\text{e}{-05}$
 (b) Second step fit with extracted scaling parameters.
@@
 Additional results, including performance on textbook-style datasets, are provided in Table 3 in the Appendix.
+4.5 Why do Diffusion models outperform AR models in data-constrained settings?
+To better understand why diffusion models are more data-efficient than autoregressive (AR) models, we conducted a series of controlled experiments aimed at isolating the core source of diffusion’s advantage.
+Our hypothesis is that masked diffusion models benefit from a form of implicit data augmentation arising from their randomized masking and denoising objective.
+Unlike AR models, which are trained exclusively on a fixed left-to-right factorization, diffusion models are exposed to a diverse set of conditional prediction tasks due to random masking patterns, enabling them to learn to generate tokens in varying orders.
+(a) Validation loss under varying attention dropout levels in AR training.
+(b) Validation loss under varying token masking levels in AR training.
+Figure 7: Impact of common data augmentations on AR models.
+Despite applying attention dropout and token masking, AR models still overfit and underperform compared to diffusion models.
+We believe this gap arises because diffusion models learn random factorizations of the joint distribution, rather than a fixed left-to-right ordering.
+To test whether this diversity could be replicated in AR models through explicit augmentation, we first experimented with standard perturbation-based techniques during AR training.
+Specifically, we applied (i) attention dropout, where a random fraction of attention weights was dropped at varying levels (25%, 50%, 75%), and (ii) token masking, where a subset of input tokens was masked by zeroing out their attention weights across all layers, while retaining the next-token prediction objective.
+As shown in Figure 7(a) and Figure 7(b), neither of these augmentations improved validation loss.
+In all cases, AR models continued to overfit and remained significantly behind diffusion models trained for longer epochs.
+Figure 8: Validation loss improves as the number of token orderings NN increases in AR training. At $N=16$, performance approaches that of diffusion models.
+We next investigated whether the key factor behind diffusion’s performance lies in its exposure to diverse token orderings.
+To evaluate this, we trained AR models with varying numbers of token orderings: $N=1$ corresponds to standard left-to-right training, while $N=k$ includes the left-to-right order along with k1k{-}1 additional random permutations of the token sequence.
+As shown in Figure 8, increasing $N$ consistently reduced validation loss and delayed overfitting.
+At $N=16$, the AR models approached the performance of diffusion models, suggesting that training over diverse orderings is indeed a major driver of diffusion’s data efficiency.
+These findings support our interpretation that diffusion models outperform AR models in low-data regimes because they are implicitly trained over a rich distribution of conditional prediction tasks.
+While AR models can partially recover this behavior via explicit random orderings, they remain fundamentally constrained by their fixed prediction structure.
+Finally, this analysis suggests a natural interpolation between the two paradigms.
+By controlling the diversity of prediction tasks—through masking or reordering—it may be possible to design hybrid models that trade off between compute efficiency (AR-like behavior) and data efficiency (diffusion-like behavior).
+Exploring this continuum offers a promising direction for future research.
+More details on how we permute the sequence can be found in 11
 Table 2: Downstream Results for the best autoregressive and diffusion trained in different data-constrainted settings.
@@
 5 Discussion
-Why do diffusion models outperform autoregressive (AR) models in low-data regimes?
-We hypothesize that the key advantage stems from the random factorizations in diffusion objective, which serves as a form of implicit data augmentation.
-Unlike AR models, which are trained on a single, fixed left-to-right ordering, masked diffusion models are exposed to a wide variety of token prediction tasks and orderings during training—including, the left-to-right order itself.
-This broader distribution over prediction tasks encourages better generalization and more effective use of each training example.
-Analogous to image-based tasks where techniques like random cropping or color jittering boost generalization  caron2021emerging ; cubuk2020randaugment ; he2022masked , we believe the corruption process in diffusion plays a similar role.
-By learning to denoise across many corruptions, the model extracts richer signal per example over time, resulting in improved data efficiency.
-We further analyze this question in Section 9 in Appendix.
 Why are autoregressive (AR) models more compute-efficient than diffusion models?
@@
 6 Limitations
-In this work, we examined two extremes of generative modeling: masked diffusion models, which learn random token orderings and are more data-efficient, and autoregressive (AR) models, which follow a fixed left-to-right order and are more compute-efficient.
+In this work, we examined two extremes of generative modeling: masked diffusion models, which learn over random condition prediction tasks and are more data-efficient, and autoregressive (AR) models, which follow a fixed left-to-right order and are more compute-efficient.
 While our results highlight a clear trade-off, this need not be binary—hybrid models that interpolate between AR and diffusion would offer a better balance.
 Although prior works have explored such hybrids (arriola2025block, ; hoogeboom2021autoregressive, ), they have not been evaluated through the lens of data-compute efficiency.
+We explore part of this question in Section 4.5, however it will be useful to study this in more detail.
 Additionally, our scaling laws are currently fit over a limited range of unique data sizes; extending them to larger regimes may improve predictive accuracy and reveal further insights.
@@
 We thank Stella Biderman for pointing out the inconsistent unit used in the Critical Compute equation.
+We thank Lucas Beyer and You Jiacheng, in suggesting the experiments in Section 4.5, that helped us explain why diffusion models would be more data-efficient.
 This work was supported in part by ONR MURI N00014-22-1-2773, ONR N00014-22-1-2096, and AFOSR FA9550-23-1-0747.

### v6

@@
 Diffusion models make better use of repeated data, achieving lower validation loss and superior downstream performance.
-We interpret this advantage as implicit data augmentation: masked diffusion exposes the model to a diverse distribution of token orderings and prediction tasks, unlike AR’s fixed left-to-right factorization.
 We find new scaling laws for diffusion models and derive a closed-form expression for the critical compute threshold at which diffusion begins to outperform AR.
-These results suggest that when data, not compute, is the bottleneck, diffusion models offer a compelling alternative to the standard AR paradigm.
+Finally, we explain why diffusion models excel in this regime: their randomized masking objective implicitly trains over a rich distribution of token orderings, acting as an implicit data augmentation that AR’s fixed left-to-right factorization lacks.
+Our results suggest that when data, not compute, is the bottleneck, diffusion models offer a compelling alternative to the standard AR paradigm.
 Our code is available at: https://diffusion-scaling.github.io.
@@
 We find the above benefits extend beyond validation loss: the best diffusion model trained in data-constrained settings consistently outperform the best autoregressive model on a range of downstream language tasks (Section 4.4).
+  6. 6.
+    Exposure to different token orderings helps explain diffusion’s data efficiency.
+By adding explicit data augmentations to AR training, we find that diffusion models’ advantage arises from their exposure to a diverse set of token orderings.
+Essentially, the randomized masking in diffusion’s objective serves as implicit data augmentation, allowing it to generalize beyond the fixed left-to-right factorization of AR models.
+(Section 4.5)
 Through detailed scaling law analysis and downstream task evaluations, we demonstrate that diffusion models make significantly better use of repeated data, achieving lower validation loss and better generalization to downstream tasks.
@@
 These scaling laws allow us to analyze performance trends and identify scenarios where diffusion models should be preferred over autoregressive ones (Section 4.3).
-Finally, in Section 4.4, we demonstrate that the superior validation loss of diffusion models indeed correlates with improved downstream task performance.
+In Section 4.4, we demonstrate that the superior validation loss of diffusion models indeed correlates with improved downstream task performance.
+Finally, in Section 4.5 we investigate the underlying cause of diffusion’s advantage in data-constrained settings, showing that its exposure to diverse token orderings enables better generalization than AR’s fixed left-to-right factorization.
 4.1 Does Diffusion Beat AR in Data-Constrained Settings?
@@
 Additional results, including performance on textbook-style datasets, are provided in Table 3 in the Appendix.
-4.5 Why do Diffusion models outperform AR models in data-constrained settings?
-To better understand why diffusion models are more data-efficient than autoregressive (AR) models, we conducted a series of controlled experiments aimed at isolating the core source of diffusion’s advantage.
-Our hypothesis is that masked diffusion models benefit from a form of implicit data augmentation arising from their randomized masking and denoising objective.
-Unlike AR models, which are trained exclusively on a fixed left-to-right factorization, diffusion models are exposed to a diverse set of conditional prediction tasks due to random masking patterns, enabling them to learn to generate tokens in varying orders.
-(a) Validation loss under varying attention dropout levels in AR training.
-(b) Validation loss under varying token masking levels in AR training.
-Figure 7: Impact of common data augmentations on AR models.
-Despite applying attention dropout and token masking, AR models still overfit and underperform compared to diffusion models.
-We believe this gap arises because diffusion models learn random factorizations of the joint distribution, rather than a fixed left-to-right ordering.
-To test whether this diversity could be replicated in AR models through explicit augmentation, we first experimented with standard perturbation-based techniques during AR training.
-Specifically, we applied (i) attention dropout, where a random fraction of attention weights was dropped at varying levels (25%, 50%, 75%), and (ii) token masking, where a subset of input tokens was masked by zeroing out their attention weights across all layers, while retaining the next-token prediction objective.
-As shown in Figure 7(a) and Figure 7(b), neither of these augmentations improved validation loss.
-In all cases, AR models continued to overfit and remained significantly behind diffusion models trained for longer epochs.
-Figure 8: Validation loss improves as the number of token orderings NN increases in AR training. At $N=16$, performance approaches that of diffusion models.
-We next investigated whether the key factor behind diffusion’s performance lies in its exposure to diverse token orderings.
-To evaluate this, we trained AR models with varying numbers of token orderings: $N=1$ corresponds to standard left-to-right training, while $N=k$ includes the left-to-right order along with k1k{-}1 additional random permutations of the token sequence.
-As shown in Figure 8, increasing $N$ consistently reduced validation loss and delayed overfitting.
-At $N=16$, the AR models approached the performance of diffusion models, suggesting that training over diverse orderings is indeed a major driver of diffusion’s data efficiency.
-These findings support our interpretation that diffusion models outperform AR models in low-data regimes because they are implicitly trained over a rich distribution of conditional prediction tasks.
-While AR models can partially recover this behavior via explicit random orderings, they remain fundamentally constrained by their fixed prediction structure.
-Finally, this analysis suggests a natural interpolation between the two paradigms.
-By controlling the diversity of prediction tasks—through masking or reordering—it may be possible to design hybrid models that trade off between compute efficiency (AR-like behavior) and data efficiency (diffusion-like behavior).
-Exploring this continuum offers a promising direction for future research.
-More details on how we permute the sequence can be found in 11
 Table 2: Downstream Results for the best autoregressive and diffusion trained in different data-constrainted settings.
@@
 Note: All values represent accuracy (%). Best results shown in bold.
+4.5 Why do Diffusion models outperform AR models in data-constrained settings?
+To better understand why diffusion models are more data-efficient than autoregressive (AR) models, we conducted a series of controlled experiments aimed at isolating the core source of diffusion’s advantage.
+Our hypothesis is that masked diffusion models benefit from a form of implicit data augmentation arising from their randomized masking and denoising objective.
+Unlike AR models, which are trained exclusively on a fixed left-to-right factorization, diffusion models are exposed to a diverse set of conditional prediction tasks due to random masking patterns, enabling them to learn to generate tokens in varying orders.
+(a) Validation loss under varying attention dropout levels in AR training.
+(b) Validation loss under varying token masking levels in AR training.
+Figure 7: Impact of common data augmentations on AR models.
+Despite applying attention dropout and token masking, AR models still overfit and underperform compared to diffusion models.
+We believe this gap arises because diffusion models learn random factorizations of the joint distribution, rather than a fixed left-to-right ordering.
+Figure 8: Validation loss improves as the number of token orderings NN increases in AR training. At $N=16$, performance approaches that of diffusion models.
+To test whether the diversity in diffusion training could be replicated in AR models via explicit augmentation, we first applied standard perturbation-based techniques during AR training.
+Specifically, we used: (i) attention dropout — randomly dropping 25%, 50%, or 75% of attention weights; and (ii) token masking — masking a subset of input tokens by zeroing their attention weights across all layers, while retaining the standard next-token prediction objective.
+As shown in Figures 7(a) and 7(b), neither approach improved validation loss.
+In all cases, AR models continued to overfit and remained far behind diffusion models trained for longer epochs.
+All AR baselines here used 140M parameters and were trained for 50 epochs; the red line in the plots marks the best diffusion model from Figure 4(a), trained for 500 epochs.
+We next investigated whether diffusion’s advantage stems from exposure to diverse token orderings.
+To test this, we trained AR models with varying numbers of orderings: $N=1$ denotes standard left-to-right training, while $N=k$ adds k1k{-}1 random permutations of the sequence order.
+All the AR models in this setting had 278M parameters and were trained for 100 epochs.
+As shown in Figure 8, increasing $N$ consistently lowered validation loss and delayed overfitting.
+At $N=16$, the 100-epoch validation loss of AR models approached that of diffusion, suggesting that diverse orderings are indeed a key driver of diffusion’s data efficiency.
+These results support our interpretation that diffusion models outperform AR models in low-data regimes because they are implicitly trained on a richer distribution of conditional prediction tasks.
+Finally, this analysis suggests a natural continuum between the two paradigms: by controlling task diversity—through masking or reordering—we could design hybrid models that interpolate between compute efficiency (AR-like) and data efficiency (diffusion-like).
+Exploring this continuum is a promising direction for future work.
+Details of our permutation process are in Section 11.
 5 Discussion

### v7

@@
 Training large language models (LLMs) on massive corpora of internet text has become the driver of recent AI breakthroughs (brown2020language, ; radford2019language, ; touvron2023llama, ).
-This progress has been fueled by scaling two core resources proportionately: compute and data  (kaplan2020scaling, ; hoffmann2022training, ).
-While compute availability has steadily grown—enabled by advances in hardware and the construction of larger data centers—the growth in high-quality data has been comparatively stagnant.
+This progress has been fueled by scaling two core resources: compute and data  (kaplan2020scaling, ; hoffmann2022training, ).
+While compute availability is steadily growing—enabled by advances in hardware and the construction of larger data centers—the growth in high-quality data is stagnating.
 Recent projections, such as those by Villalobos et.al.
-(villalobos2022will, ), estimate that the global supply of publicly available, human-generated data may be exhausted by 2028, posing a serious bottleneck to further scaling.
-This looming constraint makes it increasingly important to develop modeling strategies that are more data-efficient.
+(villalobos2022will, ), estimate that the global supply of publicly available, human-generated data may be exhausted in the coming years, posing a serious bottleneck to further scaling.
+This looming constraint makes it increasingly important to develop modeling strategies that are more sample-efficient.
 Furthermore, there are several domains, such as robotics and healthcare, where the data, not compute, is a scarce resource even to begin with.
@@
 While this modeling approach has delivered state-of-the-art performance across a range of benchmarks, it remains unclear whether it is the optimal strategy going forward.
-Recently, diffusion-based models—specifically masked diffusion models  austin2021structured ; sahoo2024simple ; lou2310discrete ; shi2024simplified ; arriola2025block —have emerged as an alternative strategy, framing text generation as an iterative masked denoising process rather than next-token prediction.
+Recently, diffusion-based models—specifically masked diffusion models  austin2021structured ; shi2024simplified ; lou2310discrete ; swerdlow2025unidisc ; arriola2025block —have emerged as an alternative strategy, framing text generation as an iterative masked denoising process rather than next-token prediction.
 At each step, the model predicts a randomly masked subset of tokens conditioned on the remaining ones, implicitly averaging over many conditional prediction orders instead of committing to one.
@@
 But a critical nuance is often overlooked: these comparisons are based entirely on single-epoch training, where each token is seen only once.
-This conflates compute efficiency with data efficiency, making it unclear whether diffusion models truly need 16× more compute—or simply 16× more data.
+This conflates compute efficiency with sample efficiency, making it unclear whether diffusion models truly need 16× more compute—or simply 16× more data.
 To resolve this ambiguity, we systematically study masked diffusion models in data-constrained settings, where repeated training on limited data is the norm rather than the exception.
@@
   6. 6.
-    Exposure to different token orderings helps explain diffusion’s data efficiency.
+    Exposure to different token orderings helps explain diffusion’s sample efficiency.
 By adding explicit data augmentations to AR training, we find that diffusion models’ advantage arises from their exposure to a diverse set of token orderings.
@@
 These results suggest that diffusion models may offer a compelling and underappreciated advantage in scenarios where data—not compute—is the primary bottleneck.
-2 Related Work
-Deep Learning in Data-Constrainted Settings.
-Deep learning progress has been largely driven by the scaling of both data and compute.
-However, recent analyses suggest we may soon face a data bottleneck that could inhibit continued advancement (villalobos2022will, ).
-In language modeling, the dominant paradigm has been autoregressive (AR) models (vaswani2017attention, ; radford2019language, ; brown2020language, ), which are typically trained for a single epoch to maximize exposure to unique tokens (hoffmann2022training, ).
-In light of looming data constraints, Muennighoff et al.(muennighoff2023scaling, ) show that AR models can still benefit from data reuse: training for up to four epochs on repeated data achieves performance nearly on par with training on fresh data, suggesting an effective strategy for improving data efficiency.
-In contrast, computer vision has long embraced multi-epoch training along with aggressive data augmentation—such as random cropping, flipping, and color jittering—to expand effective dataset size and improve generalization(shorten2019survey, ; yang2022image, ), particularly for discriminative tasks like classification and detection.
-Despite these practices, data efficiency in generative modeling remains underexplored, and the trade-offs between leading paradigms such as diffusion and AR models under constrained data regimes are still poorly understood.
-Diffusion-Based Language Models.
-[…]
 2 Method
 Our objective is to determine whether masked diffusion language models are more effective than standard autoregressive models in data-constrainted settings.
-The main difference between AR and diffusion models is the way they factorize the joint distribution of the sequence.
-Masked diffusion factorizes the joint distribution of the sequence in a random order, while AR factorizes the joint distribution of the sequence in a left-to-right order.
-To isolate the impact of this, we keep the architecture and data pipeline fixed across both families and vary only the factorization of the joint distribution.
+For studying this, we keep the core architecture and data pipeline fixed across both families.
 2.1 Preliminaries:
 Autoregressive models.
-In Autoregressive LLMs (touvron2023llama, ; radford2019language, ; brown2020language, ) we predict each token based on a growing prefix of prior tokens, defining a left-to-right factorization of the sequence probability:
+In Autoregressive LLMs (touvron2023llama, ; radford2019language, ; brown2020language, ) each token is predicted based on a growing prefix of prior tokens, defining a left-to-right factorization of the sequence probability:
 $p_{\text{AR}}(x_{1},\dots,x_{L})=\prod_{j=1}^{L}p(x_{j}\mid x_{<j}).$
@@
 Diffusion models.
-Masked diffusion language models  (austin2021structured, ; sahoo2024simple, ; nie2024scaling, ; swerdlow2025unidisc, ) treat generation as iterative denoising.
-For each training sequence $x=(x_{1},\dots,x_{L})$ we
+Masked diffusion language models  (austin2021structured, ; shi2024simplified, ; nie2024scaling, ; swerdlow2025unidisc, ) treat generation as iterative denoising.
+For each training sequence $x=(x_{1},\dots,x_{L})$ diffusion models
 1.
@@
 2. Denoise by predicting the original tokens at the masked positions with full (bidirectional) attention over $\tilde{x}$:
-$p_{\text{Diffusion}}(x\mid\tilde{x})\;=\;\prod_{i\in\mathcal{M}}p_{\theta}\!\bigl{(}x_{i}\,\mid\,\tilde{x}\bigr{)}$.
-Because the mask pattern is resampled for every example, the model is implicitly trained on a vast collection of token–ordering tasks; the standard left-to-right ordering used by AR models is just one ordering within this ensemble.
-The absence of a causal mask allows each prediction to attend to both past and future unmasked tokens, making the factorization fundamentally non-sequential.
+[…]
+Because the mask pattern is resampled for every example, the model is implicitly trained on a vast collection of token–ordering tasks.
+The absence of a causal mask allows each prediction to attend to both past and future unmasked tokens.
 2.2 Modeling Details for AR and Masked Diffusion
-Our goal is to isolate the impact of the factorization—fixed left-to-right versus random-order denoising—while keeping every other design choice constant.
-Unless noted otherwise, both model families share the same Transformer backbone (GPT-2 style with rotary positional embeddings, RoPE (su2021roformer, )) and identical training hyper-parameters across the full parameter sweep (7 M – 2 B).
+Both model families share the same Transformer backbone (GPT-2 style with rotary positional embeddings, RoPE (su2021roformer, )).
 Given a clean input sequence $x=(x_{1},\dots,x_{L})\in\mathcal{V}^{L}$, both models minimize a token-level cross-entropy loss, yet they differ in the conditioning context:
@@
 AR models predict each token conditioned on its prefix using a causal attention mask:
-$\mathcal{L}_{\text{AR}}\;=\;-\sum_{j=2}^{L}\log p_{\theta}\!\bigl{(}x_{j}\mid x_{<j}\bigr{)}.$
+$\mathcal{L}_{\text{AR}}\;=\;-\sum_{j=1}^{L}\log p_{\theta}\!\bigl(x_{j}\mid x_{<j}\bigr).$
 Masked Diffusion objective.
@@
 The loss is then
-$\mathcal{L}_{\text{Diffusion}}\;=\;-\mathbb{E}_{r}\;\mathbb{E}_{\tilde{x}\sim q_{r}}\;\frac{1}{r}\sum_{i\in\mathcal{M}}\log p_{\theta}\!\bigl{(}x_{i}\mid\tilde{x}\bigr{)}$.
+[…]
+which can be interpreted as an evidence lower bound (ELBO) on the data log-likelihood, and thus the loss provides an upper bound on the true negative log-likelihood.
 Beyond the attention mechanism and input corruption, all other variables are held constant.
@@
 In particular, we use a dynamic learning rate schedule that adapts to the number of training epochs.
-The only distinctions between AR and diffusion models in our implementation are:
-  1. 1.
-    Attention mechanism: Causal attention for AR; full self-attention for masked diffusion.
-  2. 2.
-    Prediction target: AR models predict the next token; diffusion models predict the masked tokens.
+For more detailed information on both model families please refer to related work Section in Appendix 6
 2.3 Scaling Framework in Data-Constrained Settings
-Classical scaling laws, such as those proposed by (kaplan2020scaling, ; hoffmann2022chinchilla, ), model validation loss as a function of total parameters ($N$) and training tokens ($D$), assuming all data is unique.
-These laws have been instrumental in guiding compute-optimal training of language models.
-However, this assumption becomes unrealistic as the community approaches the limits of high-quality text data available on the internet.
-To address this, Muennighoff et al.muennighoff2023scaling extend the Chinchilla framework to explicitly account for repeated data — a common necessity in data-constrained regimes.
+Classical scaling laws, such as those proposed by (kaplan2020scaling, ; hoffmann2022chinchilla, ), model validation loss as a function of total parameters ($N$) and training tokens ($D$), assuming all data is unique that is single epoch regime.
+Muennighoff et al.muennighoff2023scaling extend the Chinchilla framework to explicitly account for repeated data — a common necessity in data-constrained regimes.
 They show that repeating training data beyond a few epochs yields diminishing returns and propose a new scaling law that incorporates the decaying utility of repeated tokens.
 We briefly outline their formulation below.
-Definitions:
-  • $U$: number of unique tokens available for training,
-  • $E$: number of epochs (i.e., how many times each unique token is reused),
-  • $D=U\cdot E$: total number of tokens seen by the model.
+Let $U$ denote the number of unique tokens, $E$ the number of epochs (how many times each unique token is reused), and $D=U\cdot E$ the total number of tokens seen during training.
 To model diminishing returns from repeated data, Muennighoff et al.
@@
 This defines a natural saturation point on returns: even infinite compute yields no additional effective data beyond this limit.
-A symmetric formulation is applied to model parameters for mathetmatical convenience which is used to define $N^{\prime}$.
+A symmetric formulation is applied to model parameters for mathematical convenience which is used to define $N^{\prime}$.
 Finally, a modified Chinchilla-style loss function incorporates these effective quantities $N^{\prime}$ and $D^{\prime}$:
@@
 This may provide a slight advantage to autoregressive models, as these hyperparameters were originally tuned for that family.
-For all models, we use the following hyperparameters: batch size of 256 sequences, AdamW optimizer with $\beta_{1}{=}0.9$, $\beta_{2}{=}0.95$, $\epsilon{=}10^{-8}$, a learning rate schedule with peak 2e-4, minimum 2e-5, 1% warm-up, cosine decay, weight decay 0.1, and gradient clipping of 1.0.
+For details on hyperparameters please refer to Section 8 in Appendix.
 3 Experiments
@@
 As a result, they do not isolate whether diffusion’s 16x inefficiency stems from needing more total compute—or more unique data.
-In other words: is diffusion limited by compute efficiency or by data efficiency?
+In other words: is diffusion limited by compute efficiency or by sample efficiency?
 To answer this, we systematically study diffusion models in data-constrained settings, where the total amount of unique data is fixed and models are trained for many epochs, reusing the same data.
 Unlike prior work, our evaluation explicitly decouples model scaling from data reuse, allowing us to disentangle the effects of compute and data.
-We train a large suite of AR and diffusion models across three unique data regimes—25M, 50M, and 100M tokens—and a wide range of training compute budgets.
 In Figure 1, we report empirical validation loss as a function of training FLOPs for the 50M and 100M regimes; results for the 25M setting are shown in Appendix Figure 9.
@@
 However, as training is extended to hundreds of epochs, diffusion models continue to improve and eventually achieve a lower validation loss (3.55) than the best AR models (3.71).
-While AR models begin to overfit as training progresses, diffusion models show no signs of overfitting within our budget.
 (a) Autoregressive contour: validation loss over epochs and model sizes.
@@
 This corresponds to a 67% reduction in loss for diffusion models compared to just 48% for AR models, highlighting their superior ability to leverage repeated data.
-These results underscore that diffusion models require significantly more training—both in epochs and compute—to realize their advantages in data-constrained settings.
 3.2 Fitting Data-Constrained Scaling Laws
+Figure 3: Decay rate of data value under repetition: left shows diffusion, middle AR, and right the average decay rate for both.
+Points are empirical results (darker color = higher FLOPs, lighter color = lower FLOPs; each line = fixed compute), we find that fitted curves (represented as lines) closely match the empirical points, indicating our scaling laws are representative.
+The decay rate of value for repeated data is lower for diffusion, reflecting its greater robustness to repeating.
+(a) AR Training Curves.
+(b) Diffusion Training Curves.
+Figure 4: Training curves for different epoch counts, all with using the same total compute.
+Each curve shows a different tradeoff between unique data and repetition.
+For AR models, validation loss rises with more epochs (overfitting), while for diffusion models, the curves are nearly unchanged, showing much greater robustness to data repetition.
+Figure 5: Predicted validation loss for AR (left) and Diffusion models (right) under compute-optimal settings, extrapolated to larger compute budgets.
+Dotted lines show the hypothetical case where repeated data equals new data.
+For AR, this holds up to $\approx$4 epochs; for diffusion, up to $\approx$100 epochs—showing diffusion’s greater robustness to data repetition.
+Note that loss values between AR and diffusion are not directly comparable, as they’re extrapolated from scaling laws with different data-entropy terms ($E_{0}$).
+In Section 3.5, we ignore this factor during comparison.
 To gain deeper insight into the trade-offs between diffusion and autoregressive models in data-constrained settings, we fit scaling laws to both model families across single-epoch and multi-epoch regimes, as described in Section 2.3.
@@
 A higher $R_{D}^{*}$ implies that a model can benefit from many more repeated exposures before saturating.
-This suggests that diffusion models continue to improve across hundreds of epochs, while AR models quickly saturate—highlighting the superior data efficiency of diffusion models in data-constrained regimes.
+This suggests that diffusion models continue to improve across hundreds of epochs, while AR models quickly saturate—highlighting the superior sample efficiency of diffusion models in data-constrained regimes.
 Table 1: Fitting metrics of the scaling law model for Diffusion and AR. Diffusion and AR achieve a strong fit across both phases.
@@
 In contrast, diffusion models continue to match the unique-data curve for up to 100 epochs, indicating a far greater capacity to benefit from repeated data in data-constrained regimes.
-Figure 3: Decay rate of data value under repetition: left shows diffusion, middle AR, and right the average decay rate for both.
-Points are empirical results (darker color = higher FLOPs, lighter color = lower FLOPs; each line = fixed compute), we find that fitted curves (represented as lines) closely match the empirical points, indicating our scaling laws are representative.
-The decay rate of value for repeated data is lower for diffusion, reflecting its greater robustness to repeating.
-(a) AR Training Curves.
-(b) Diffusion Training Curves.
-Figure 4: Training curves for different epoch counts, all with using the same total compute.
-Each curve shows a different tradeoff between unique data and repetition.
-For AR models, validation loss rises with more epochs (overfitting), while for diffusion models, the curves are nearly unchanged, showing much greater robustness to data repetition.
-Figure 5: Predicted validation loss for AR models (left) and Diffusion models (right) under compute-optimal settings, extrapolated to larger compute budgets.
-Dotted lines indicate the hypothetical case where repeated data is as valuable as new data.
-For AR, this holds up to about 4 epochs; for diffusion, up to 100 epochs, showing that diffusion models are much more robust to data repetition.
 3.3 When to Use Diffusion over AR?
-A key question for practitioners is: when should diffusion be preferred over autoregressive models (AR)?
-To answer this, we compare the fitted data-constrained scaling laws for both model families (§3.3).
-We define the validation loss gap between diffusion and AR as:
-$\Delta\mathcal{L}(C,U)=\mathcal{L}_{\text{Diffusion}}(C,U)-\mathcal{L}_{\text{AR}}(C,U),$
-where $C$ is total training compute and $U$ is the number of unique tokens.
-Positive values favor AR; negative values favor diffusion.
-The critical compute $C_{\mathrm{crit}}(U)$ is the point where the models perform equally:
-$\Delta\mathcal{L}(C_{\mathrm{crit}},U)=0.$
-Figure 6(a) shows a heatmap of $\Delta\mathcal{L}$ over compute and data.
-Red regions indicate regimes where diffusion outperforms AR ($\Delta\mathcal{L}<0$), while blue regions favor AR.
-As expected, AR performs better in low-compute settings due to its efficient per-step learning.
-However, diffusion models begin to outperform AR at higher compute, especially when data is limited and repeated.
-Figure 6(b) plots the critical compute frontier $C_{\mathrm{crit}}(U)$—the compute required for diffusion to match AR at a given unique token count $U$. This frontier follows a power law:
-$C_{\mathrm{crit}}(U)\propto U^{2.174}.$
-The linear fit in log-log space is:
-$\log_{10}(U)=0.460\cdot\log_{10}(C)-1.050,\quad\text{so}\quad C_{\text{crit}}(U)=2.12\times 10^{1.956}\cdot U^{2.174}.$
-The green dashed line shows the fitted curve, and the blue crosses represent empirical crossover points—where diffusion matches AR performance in experiments.
-These points align closely with the predicted frontier, confirming our fitted equation’s accuracy.
 (a) Loss Gap Heatmap.
@@
 Right: The critical compute curve defining the compute threshold needed for diffusion to match autoregressive models at a given unique token count.
+A key question for practitioners is: when should diffusion be preferred over autoregressive models (AR)?
+To answer this, we compare the fitted data-constrained scaling laws for both model families (§2.3).
+We define the validation loss gap between diffusion and AR as:
+$\Delta\mathcal{L}(C,U)=\mathcal{L}_{\text{Diffusion}}(C,U)-\mathcal{L}_{\text{AR}}(C,U),$
+where $C$ is total training compute and $U$ is the number of unique tokens.
+Positive values favor AR; negative values favor diffusion.
+The critical compute $C_{\mathrm{crit}}(U)$ is the point where the models perform equally: $\Delta\mathcal{L}(C_{\mathrm{crit}},U)=0.$
+Figure 6(a) shows a heatmap of $\Delta\mathcal{L}$ over compute and data.
+Red regions indicate regimes where diffusion outperforms AR ($\Delta\mathcal{L}<0$), while blue regions favor AR.
+As expected, AR performs better in low-compute settings due to its efficient per-step learning.
+However, diffusion models begin to outperform AR at higher compute, especially when data is limited and repeated.
+Figure 6(b) plots the critical compute frontier $C_{\mathrm{crit}}(U)$—the compute required for diffusion to match AR at a given unique token count $U$.
+This frontier follows a power law: $C_{\mathrm{crit}}(U)\propto U^{2.174}.$ The linear fit in log-log space is:
+$\log_{10}(U)=0.460\cdot\log_{10}(C)-1.050,\quad\text{so}\quad C_{\text{crit}}(U)=2.12\times 10^{1.956}\cdot U^{2.174}.$
+The dark green dashed line shows the fitted curve, and the blue stars represent empirical crossover points—where diffusion matches AR performance in experiments.
+These points align closely with the predicted frontier, confirming our fitted equation’s accuracy.
 3.4 Downstream Results
-We evaluate the best-performing diffusion and autoregressive (AR) models on several downstream benchmarks to assess whether the gains in validation loss translate to practical improvements in generalization.
-Motivated by the critical compute threshold equation identified in Section 4.3, we scale the training data to 500M unique tokens and train a 2.3B parameter diffusion model using the compute budget predicted by the critical compute limit.
-We train the model for 130 epochs, during which we observe no signs of convergence.
-We terminate the training due to compute constraints.
+Benchmarks Random Baseline AR AR (Flop matched) Diffusion
+ARC-Easy clark2018think 25.00 35.63 35.35 37.84
+BoolQ clark2019boolq 50.00 46.00 38.23 49.38
+COPA roemmele2011choice 50.00 56.33 54.00 59.00
+HellaSwag zellers2019hellaswag 25.00 27.37 29.03 30.24
+PiQA 50.00 60.94 61.64 60.72
+RACE lai2017race 25.00 25.28 24.88 28.96
+WinoGrande XL sakaguchi2021winogrande 50.00 48.87 49.41 50.97
+SciQ SciQ 25.00 58.05 50.50 68.67
+Lambada paperno2016lambada 00.00 10.91 5.53 15.19
+Note: All values represent accuracy (%). Best results shown in bold.
+Table 2: Downstream results for the best-performing (as per validation loss) and flop-matched autoregressive (AR) and diffusion models trained with 100M unique tokens.
+Random baselines are reported for reference.
+We evaluate the best-performing diffusion and autoregressive (AR) models, selected based on their validation loss, across several downstream benchmarks to examine whether lower validation loss translates to improved generalization.
+Since AR models tend to overfit much earlier, we additionally evaluate flop-matched overfitted AR models trained for the same number of epochs as their diffusion counterparts.
+On a few set of benchmarks, these overfitted AR models slightly outperform their best-validation variants; however, across almost all tasks, diffusion models consistently achieve the highest downstream performance.
+Additional results, with 500M unique tokens and other downstream datasets, are provided in Table 3 and Table 4 in the Appendix.
 Across a diverse set of tasks and data scales, diffusion models consistently outperform their AR counterparts.
-This validates our findings in Section 4.3, also confirms that the data efficiency gains observed in validation loss translate into stronger downstream performance.
-Additional results, including performance on textbook-style datasets, are provided in Table 3 in the Appendix.
-Table 2: Downstream Results for the best autoregressive and diffusion trained in different data-constrainted settings.
-We report the results for the models with the best validation loss in 100M and 500M unique data regime.To better understand the difficulty of each benchmark we also report the accuracy of random baseline
-Benchmarks Random Baseline 100M unique tokens 500M unique tokens
-AR Diffusion AR Diffusion
-ARC-Easy clark2018think 25.00 35.63 37.84 43.79 45.95
-BoolQ clark2019boolq 50.00 46.00 49.38 51.87 55.26
-COPA roemmele2011choice 50.00 56.33 59.00 67.00 64.83
-HellaSwag zellers2019hellaswag 25.00 27.37 30.24 32.28 35.33
-PiQA 50.00 60.94 60.72 65.71 65.61
-RACE lai2017race 25.00 25.28 28.96 28.28 31.44
-WinoGrande XL sakaguchi2021winogrande 50.00 48.87 50.97 50.61 51.51
-SciQ SciQ 25.00 58.05 68.67 67.82 79.13
-Lambada paperno2016lambada 00.00 10.91 15.19 15.07 22.30
-Note: All values represent accuracy (%). Best results shown in bold.
 3.5 Why do Diffusion models outperform AR models in data-constrained settings?
-To better understand why diffusion models are more data-efficient than autoregressive (AR) models, we conducted a series of controlled experiments aimed at isolating the core source of diffusion’s advantage.
-Our hypothesis is that masked diffusion models benefit from a form of implicit data augmentation arising from their randomized masking and denoising objective.
-Unlike AR models, which are trained exclusively on a fixed left-to-right factorization, diffusion models are exposed to a diverse set of conditional prediction tasks due to random masking patterns, enabling them to learn to generate tokens in varying orders.
-(a) Validation loss under varying attention dropout levels in AR training.
-(b) Validation loss under varying token masking levels in AR training.
-Figure 7: Impact of common data augmentations on AR models.
-Despite applying attention dropout and token masking, AR models still overfit and underperform compared to diffusion models.
-We believe this gap arises because diffusion models learn random factorizations of the joint distribution, rather than a fixed left-to-right ordering.
+To better understand why diffusion models are more sample-efficient than autoregressive (AR) models, we conducted a series of controlled experiments aimed at isolating the core source of diffusion’s advantage.
 Figure 7: Validation loss improves as the number of token orderings NN increases in AR training. At $N=16$, performance approaches that of diffusion models.
-To test whether the diversity in diffusion training could be replicated in AR models via explicit augmentation, we first applied standard perturbation-based techniques during AR training.
+We first applied standard perturbation-based techniques during AR training.
 Specifically, we used: (i) attention dropout — randomly dropping 25%, 50%, or 75% of attention weights; and (ii) token masking — masking a subset of input tokens by zeroing their attention weights across all layers, while retaining the standard next-token prediction objective.
-As shown in Figures 7(a) and 7(b), neither approach improved validation loss.
+As shown in Figures 8(a) and 8(b) in Appendix, neither approach improved validation loss.
 In all cases, AR models continued to overfit and remained far behind diffusion models trained for longer epochs.
-All AR baselines here used 140M parameters and were trained for 50 epochs; the red line in the plots marks the best diffusion model from Figure 4(a), trained for 500 epochs.
+All AR baselines here used 140M parameters and were trained for 50 epochs; the red line in the plots marks the best diffusion model from Figure 2(b), trained for 500 epochs.
 We next investigated whether diffusion’s advantage stems from exposure to diverse token orderings.
 To test this, we trained AR models with varying numbers of orderings: $N=1$ denotes standard left-to-right training, while $N=k$ adds k1k{-}1 random permutations of the sequence order.
-All the AR models in this setting had 278M parameters and were trained for 100 epochs.
+All permutations were fixed prior to training, and each training batch randomly samples from these orderings.
+All AR models in this setting used 278M parameters and were trained for 100 epochs.
 As shown in Figure 7, increasing $N$ consistently lowered validation loss and delayed overfitting.
-At $N=16$, the 100-epoch validation loss of AR models approached that of diffusion, suggesting that diverse orderings are indeed a key driver of diffusion’s data efficiency.
+At $N=16$, the 100-epoch validation loss of AR models approached that of diffusion, suggesting that diverse orderings are indeed a key driver of diffusion’s sample efficiency.
 These results support our interpretation that diffusion models outperform AR models in low-data regimes because they are implicitly trained on a richer distribution of conditional prediction tasks.
-Finally, this analysis suggests a natural continuum between the two paradigms: by controlling task diversity—through masking or reordering—we could design hybrid models that interpolate between compute efficiency (AR-like) and data efficiency (diffusion-like).
+Finally, this analysis suggests a natural continuum between the two paradigms: by controlling task diversity—through masking or reordering—we could design hybrid models that interpolate between compute efficiency (AR-like) and sample efficiency (diffusion-like).
 Exploring this continuum is a promising direction for future work.
 Details of our permutation process are in Section 12.
-5 Discussion
-Why are autoregressive (AR) models more compute-efficient than diffusion models?
-We hypothesize two main contributing factors.
-(i) Order specialization: AR models are trained with a fixed left-to-right factorization, so every gradient update reinforces the same prediction task, allowing them to specialize effectively.
-In contrast, diffusion models must generalize across many random token orderings, which hinders specialization.
-(ii) Stronger supervision per update: In AR training, every token in a training sequence serves as a supervised target, and the causal structure enables dense gradient updates, resulting in stable, low-variance learning.
-Diffusion models, however, compute loss only on a subset of masked tokens, making supervision sparser per sequence, even though gradients propagate through the entire input.
-As a result, each update carries less direct learning signal.
-Arriola et al.
-(arriola2025block, ) show that tuning the masking schedule can help reduce gradient variance and improve training compute efficiency.
-6 Limitations
-In this work, we examined two extremes of generative modeling: masked diffusion models, which learn over random condition prediction tasks and are more data-efficient, and autoregressive (AR) models, which follow a fixed left-to-right order and are more compute-efficient.
-While our results highlight a clear trade-off, this need not be binary—hybrid models that interpolate between AR and diffusion would offer a better balance.
-Although prior works have explored such hybrids (arriola2025block, ; hoogeboom2021autoregressive, ), they have not been evaluated through the lens of data-compute efficiency.
-We explore part of this question in Section 4.5, however it will be useful to study this in more detail.
-Additionally, our scaling laws are currently fit over a limited range of unique data sizes; extending them to larger regimes may improve predictive accuracy and reveal further insights.
 4 Conclusion
-As the availability of high-quality data plateaus, improving data efficiency becomes essential for scaling deep learning.
+As the availability of high-quality data plateaus, improving sample efficiency becomes essential for scaling deep learning.
 In this work, we show that masked diffusion models consistently outperform autoregressive (AR) models in data-constrained regimes — when training involves repeated passes over a limited dataset.
@@
 We thank Stella Biderman for pointing out the inconsistent unit used in the Critical Compute equation.
-We thank Lucas Beyer and You Jiacheng, in suggesting the experiments in Section 4.5, that helped us explain why diffusion models would be more data-efficient.
+We thank Lucas Beyer and You Jiacheng, in suggesting the experiments in Section 3.5, that helped us explain why diffusion models would be more sample-efficient.
 This work was supported in part by ONR MURI N00014-22-1-2773, ONR N00014-22-1-2096, and AFOSR FA9550-23-1-0747.
+The results and models presented in this work also used compute resources from the National AI Research Resource Pilot, with support from NVIDIA, including NVIDIA’s DGX Cloud product and the NVIDIA AI Enterprise Software Platform.
+Appendix
+6 Related Work
+Deep Learning in Data-Constrainted Settings.
+Deep learning progress has been largely driven by the scaling of both data and compute.
+However, recent analyses suggest we may soon face a data bottleneck that could inhibit continued advancement [42].
+In language modeling, the dominant paradigm has been autoregressive (AR) models [41, 28, 4], which are typically trained for a single epoch to maximize exposure to unique tokens [11].
+In light of looming data constraints, Muennighoff et al.[21] show that AR models can still benefit from data reuse: training for up to four epochs on repeated data achieves performance nearly on par with training on fresh data, suggesting an effective strategy for improving data efficiency.
+In contrast, computer vision has long embraced multi-epoch training along with aggressive data augmentation—such as random cropping, flipping, and color jittering—to expand effective dataset size and improve generalization[36, 43], particularly for discriminative tasks like classification and detection.
+Despite these practices, data efficiency in generative modeling remains underexplored, and the trade-offs between leading paradigms such as diffusion and AR models under constrained data regimes are still poorly understood.
+Diffusion-Based Language Models.
+[…]
+7 Additional Results
+(a) Validation loss under varying attention dropout levels in AR training.
+(b) Validation loss under varying token masking levels in AR training.
+Figure 8: Impact of common data augmentations on AR models.
+Despite applying attention dropout and token masking, AR models still overfit and underperform compared to diffusion models.
+We believe this gap arises because diffusion models learn random factorizations of the joint distribution, rather than a fixed left-to-right ordering.
+In Table 3, we report downstream results using 500M unique tokens.
+Guided by the critical compute threshold derived in Section 3.3, we scale training to 500M unique tokens and train a 2.3B-parameter diffusion model under the predicted compute budget.
+The model is trained for 130 epochs, after which we observe no signs of convergence and terminate due to compute limitations.
+For the autoregressive (AR) model, we report results from the checkpoint that achieves the best validation loss before overfitting..
+Figure 8(a) and Figure 8(b) show that standard perturbation-based techniques, such as attention dropout (25–75%) and token masking, fail to improve validation loss.
+Across all settings, AR models continue to overfit and lag behind diffusion models trained for longer epochs.
+All AR baselines use 140M parameters and are trained for 50 epochs, whereas the red line denotes the best diffusion model from Figure 2(b), trained for 500 epochs.
+Table 4 reports the negative log-likelihood (NLL; lower is better) on four diverse corpora: OSCAR [25], TinyStories[8], WikiText [20], and IndustryCorpus2 EN Sub [35].
+These datasets span open-domain, narrative, encyclopedic, and industry-specific text.
+Figure 9 shows validation loss versus training FLOPs for autoregressive (AR) and masked diffusion models under data-constrained settings.
+This figure extends Figure 1 in the main paper, by including empirical validation loss for the 25M unique token regimes.
+Table 3: Downstream results for the best-performing (as per validation loss) autoregressive and diffusion models trained with 500M unique tokens. We also report the random baseline for reference.
+Benchmarks Random Baseline AR Diffusion
+ARC-Easy [6] 25.00 43.79 45.95
+BoolQ [5] 50.00 51.87 55.26
+COPA [30] 50.00 67.00 64.83
+HellaSwag [45] 25.00 32.28 35.33
+PiQA 50.00 65.71 65.61
+RACE [16] 25.00 28.28 31.44
+WinoGrande XL [32] 50.00 50.61 51.51
+SciQ [14] 25.00 67.82 79.13
+Lambada [27] 00.00 15.07 22.30
+Note: All values represent accuracy (%). Best results shown in bold.
+Table 4: Downstream NLL of best diffusion and AR models at 100M unique data points.
+Model Type Flops OSCAR TinyStories WikiText IndustryCorpus2
+Best ARM 4.32e18 3.98 2.96 4.94 / 4.96 3.58
+Best MDM 1.24e20 3.83 2.93 4.50 / 4.52 3.44
+Figure 9: Pareto frontier of validation loss (negative log-likelihood) versus training FLOPs for autoregressive (AR) and diffusion models under data-constrained settings, on three different unique data settings 25M, 50M and 100M.
+8 Hyperparameter details
+We use the following hyperparameters: batch size of 256 sequences, AdamW optimizer with $\beta_{1}{=}0.9$, $\beta_{2}{=}0.95$, $\epsilon{=}10^{-8}$, a learning rate schedule with peak 2e-4, minimum 2e-5, 1% warm-up, cosine decay, weight decay 0.1, and gradient clipping of 1.0.
+9 Model Architecture
+We adopt the Megatron-DeepSpeed framework as the foundation of our implementation, upon which we build our training and evaluation setup for the masked Diffusion Model.
+Similar to the “extended version of the architectures” proposed in [22], our model adheres to the general transformer design while introducing several architectural modifications to better align with modern LLM practices.
+Specifically, we replace absolute positional embeddings with Rotary Positional Embeddings (RoPE) [37], which improve extrapolation to longer contexts and reduce parameter count.
+Furthermore, we adopt the SwiGLU activation function in the MLP blocks, which has been shown to outperform standard GELU or ReLU in both convergence and downstream performance [33].
+To further simplify the architecture and enhance training stability, we substitute standard LayerNorm with RMSNorm and eliminate all bias terms.
+These design choices are consistent with [3, 40].
+To preserve the original MLP capacity while aligning with hardware-friendly parameter sizes, we compute the feed-forward hidden size $h_{f}$ as:
+$h_{f}=\left\lfloor\frac{8\cdot d_{\text{model}}}{3\cdot 64}\right\rfloor\cdot 64$
+This rounding scheme ensures that the FFN hidden size remains divisible by 64 while closely matching the effective dimensionality used in SwiGLU layers.
+We slightly modify the parameter count estimation formula from the original:
+$P=12lh^{2}\left(1+\frac{13}{12h}+\frac{V+s}{12lh}\right)$
+to better reflect our revised architecture.
+The original formula can be decomposed into: $4lh^{2}$ (attention), $8lh^{2}$ (MLP), $13lh$ (LayerNorm and biases), and $(V+s)h$ (token and positional embeddings).
+After applying our architectural adjustments—namely, using a SwiGLU-based MLP of dimension $h_{f}$, switching to RoPE (eliminating $sh$), and removing bias terms—we arrive at the revised formula:
+$P=4lh^{2}+3lh\cdot h_{f}+6lh+Vh$
+Table 5 presents all model configurations used in our experiments along with their parameter counts.
+Table 5: Model Architectures
+Name param (M) d_model origin_ffw_size ffw_size kv_size n_heads n_layers
+7 7.0 128 512 320 32 4 3
+14 13.6 224 896 576 32 7 4
+20 19.5 288 1152 768 32 7 5
+35 36.6 448 1792 1152 32 7 6
+44 50.7 512 2048 1344 64 8 8
+57 64.8 576 2304 1536 64 9 9
+74 80.5 640 2560 1664 64 10 10
+90 95.0 640 2560 1664 64 10 13
+106 109.6 640 2560 1664 64 10 16
+117 123.6 768 3072 2048 64 12 12
+140 144.8 768 3072 2048 64 12 15
+163 166.1 768 3072 2048 64 12 18
+175 179.2 896 3584 2368 64 14 14
+196 198.3 896 3584 2368 64 14 16
+217 217.5 896 3584 2368 64 14 18
+251 250.8 1024 4096 2688 64 16 16
+278 275.7 1024 4096 2688 64 16 18
+306 300.6 1024 4096 2688 64 16 20
+425 416.9 1280 5120 3392 128 10 18
+489 475.6 1280 5120 3392 128 10 21
+509 495.9 1408 5632 3712 128 11 18
+552 534.4 1280 5120 3392 128 10 24
+587 566.7 1408 5632 3712 128 11 21
+632 615.3 1536 6144 4096 128 12 19
+664 637.6 1408 5632 3712 128 11 24
+724 700.3 1536 6144 4096 128 12 22
+816 785.2 1536 6144 4096 128 12 25
+893 856.4 1792 7168 4736 128 14 20
+1018 971.3 1792 7168 4736 128 14 23
+1143 1086.3 1792 7168 4736 128 14 26
+1266 1207.6 2048 8192 5440 128 16 22
+1424 1353.6 2176 8704 5760 128 17 22
+1429 1358.2 2048 8192 5440 128 16 25
+1593 1508.9 2048 8192 5440 128 16 28
+1609 1523.2 2176 8704 5760 128 17 25
+1731 1644.9 2304 9216 6144 128 18 24
+1794 1692.9 2176 8704 5760 128 17 28
+2007 1899.8 2304 9216 6144 128 18 28
+2283 2154.7 2304 9216 6144 128 18 32
+2298 2165.3 2560 10240 6784 128 20 26
+2639 2478.6 2560 10240 6784 128 20 30
+2980 2791.9 2560 10240 6784 128 20 34
+3530 3257.0 2688 10752 7168 128 21 36
+3802 3561.3 2816 11264 7488 128 22 36
+4084 3879.2 2944 11776 7808 128 23 36
+4516 4231.9 3072 12288 8192 128 24 36
+6796 6337.4 3584 14336 9536 128 28 40
+9293 8640.6 4096 16384 10880 128 32 42
+11452 10889.0 4352 17408 11584 128 32 47
+12295 11444.2 4608 18432 12288 128 36 44
+12569 12208.7 4608 18432 12288 128 32 47
+13735 13560.0 4864 19456 12928 128 32 47
+14940 14905.3 4992 19968 13312 128 32 49
+16183 15028.3 5120 20480 13632 128 40 47
+10 Discussion
+Why are autoregressive (AR) models more compute-efficient than diffusion models?
+We hypothesize two main contributing factors.
+(i) Order specialization: AR models are trained with a fixed left-to-right factorization, so every gradient update reinforces the same prediction task, allowing them to specialize effectively.
+In contrast, diffusion models must generalize across many random token orderings, which hinders specialization.
+(ii) Stronger supervision per update: In AR training, every token in a training sequence serves as a supervised target, and the causal structure enables dense gradient updates, resulting in stable, low-variance learning.
+Diffusion models, however, compute loss only on a subset of masked tokens, making supervision sparser per sequence, even though gradients propagate through the entire input.
+As a result, each update carries less direct learning signal.
+Arriola et al.
+[1] show that tuning the masking schedule can help reduce gradient variance and improve training compute efficiency.
+11 Limitations
+In this work, we examined two extremes of generative modeling: masked diffusion models, which learn over random condition prediction tasks and are more data-efficient, and autoregressive (AR) models, which follow a fixed left-to-right order and are more compute-efficient.
+While our results highlight a clear trade-off, this need not be binary—hybrid models that interpolate between AR and diffusion would offer a better balance.
+Although prior works have explored such hybrids [1, 13], they have not been evaluated through the lens of data-compute efficiency.
+We explore part of this question in Section 3.5, however it will be useful to study this in more detail.
+Additionally, our scaling laws are currently fit over a limited range of unique data sizes; extending them to larger regimes may improve predictive accuracy and reveal further insights.
+12 Order Permutation Details
+In this experiment, we train autoregressive models using different token orderings.
+We do not introduce target positional embeddings as done in works such as RAR  [44, 26].
+We evaluated the trained models using left-to-right ordering.
+We define the perturbations in the token ordering by adding varying levels of noise to the left-to-right ordering.
+Specifically, we generate a list of N orderings, where the first order is the standard left-to-right (l2r) order.
+Subsequent permutations are created by adding Gaussian noise to the left-to-right position ids, with the standard deviation of the noise directly proportional to the permutation’s index.
+This method allows us to create a spectrum of orderings, from the standard l2r order to more heavily permuted sequences, as detailed in Algorithm 1.
+During training, we apply these predefined orders to the input sequences.
+For each sequence in a batch, we randomly sample a permutation from our predefined list.
+This process is summarized in Algorithm 2 and further detailed below:
+For each sequence, the first token is kept fixed.
+This ensures that the position ID 0 is always assigned to the first token, providing a soft absolute positional anchor for the sequence when using RoPE[38].
+Under RoPE, attention depends only on relative position offsets rather than absolute information, i.e.
+$\langle R(i)q,R(j)k\rangle=qR(i-j)k$.
+Therefore, fixing position 0 on the first token keeps the control anchor unrotated $R(0)=I$ and removes global sequence-wise phase shifts induced by permutations, which stabilized the optimization and reduced variance under permutation augmentation.
+As an example, suppose that the number of predicted tokens is $T$ (e.g.
+$T=2048$ in our default setting) and the total input length is $L=T+1$ including the label shift.
+Only the indices in [1:T][1{:}T] are shuffled and assigned position IDs from $\{1,\dots,T\}$.
+For instance, with $T=6$ and a permutation $\pi=[2,0,1,4,5,3]$, the resulting token and label orders are:
+tokens: [ $\quad[\,0,\,3,\,1,\,2,\,5,\,6\,],$
+labels: [ $\quad[\,3,\,1,\,2,\,5,\,6,\,4\,].$
+Algorithm 1 Generating a Random Order List with Predefined Permutations
+Input: Sequence length $L$, number of orders $N$, random seed $s$
+Output: Order list $\mathcal{O}$ of $N$ orderings
+1: Initialize order list $\mathcal{O}\leftarrow[]$
+2: Append raster order: $\mathcal{O}\leftarrow\mathcal{O}\cup\{[0,1,\dots,L-1]\}$
+3:for $i=1$ to N1N-1 do
+4:  $b\leftarrow[0,1,\dots,L-1]$ {base raster order}
+5:  $\epsilon\sim\mathcal{N}(0,i^{2}I)$ {add Gaussian noise with scale ii}
+6:  $s\leftarrow b+\epsilon$ {perturbed scores}
+7:  $\pi\leftarrow\text{argsort}(s)$ {permutation order}
+8:  $\mathcal{O}\leftarrow\mathcal{O}\cup\{\pi\}$
+9:end for
+10:
+11:return $\mathcal{O}$
+Algorithm 2 Shuffling Tokens Using Predefined Order Lists
+Input: Token matrix $\texttt{tokens}\in\mathbb{Z}^{B\times L+1}$ (including last label), order list $\mathcal{O}$ of $K$ permutations
+Output: Shuffled tokens and position IDs
+1: Let $B\leftarrow$ number of sequences in batch
+2: Let $L\leftarrow$ sequence length
+3: Initialize $\texttt{position\_ids}\leftarrow\mathbf{0}^{B\times L}$
+4: Sample index vector $I\sim\text{Uniform}(\{0,\dots,K{-}1\})^{B}$ {select random order for each sequence}
+5:for $i=1$ to $B$ do
+6:  $\pi\leftarrow\mathcal{O}[I_{i}]$ {retrieve $i$-th random order}
+7:  $\texttt{tokens}[i,1{:}]\leftarrow\texttt{tokens}[i,1{:}][\pi]$ {shuffle tokens except first token}
+8:  $\pi\leftarrow\pi+1$ {shift positions by 1 to reserve position 0}
+9:  $\texttt{position\_ids}[i,1{:}]\leftarrow\pi[0{:}L{-}1]$ {assign shifted positions}
+10:end for
+11:
+12:return tokens, position_ids
+13 Broader Impacts
+Our findings suggest that masked diffusion models are more sample-efficient language modeling, which is especially valuable as high-quality textual data becomes scarce.
+This could benefit low-resource languages, privacy-sensitive domains, and scientific fields where data is limited or costly.
+By reducing dependence on massive proprietary datasets, diffusion models may help democratize access to LLM development.
+Nonetheless, the broader deployment of generative models raises concerns around misuse and misinformation, underscoring the need for responsible research and deployment practices.
+NeurIPS Paper Checklist
+  1. 1.
+    Claims
+  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
+  3. Answer: [Yes]
+  4.
+Justification: The abstract and introduction clearly state the goal of comparing diffusion and autoregressive models in data-constrained regimes.
+The main claims—that diffusion models are more sample-efficient under repeated exposures and that new scaling laws established—are consistent with the theoretical analysis and experimental results presented throughout the paper.
+  5. Guidelines:
+    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
+    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations.
+A No or NA answer to this question will not be perceived well by the reviewers.
+    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
+    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
+  6. 2.
+    Limitations
+  7. Question: Does the paper discuss the limitations of the work performed by the authors?
+  8. Answer: [Yes]
+  9. Justification: The limitations of the work, such as the restricted range of model sizes and datasets used in scaling law fitting, are explicitly discussed in the paper.
+  10. Guidelines:
+    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
+    • The authors are encouraged to create a separate "Limitations" section in their paper.
+    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally).
+The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
+    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs.
+In general, empirical results often depend on implicit assumptions, which should be articulated.
+    • The authors should reflect on the factors that influence the performance of the approach.
+For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting.
+Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
+    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
+    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
+    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper.
+The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community.
+Reviewers will be specifically instructed to not penalize honesty concerning limitations.
+  11. 3.
+    Theory assumptions and proofs
+  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
+  13. Answer: [N/A]
+  14.
+Justification: The paper does not include formal theoretical results or theorems requiring assumptions or proofs.
+While empirical scaling laws are fitted, they are based on empirical loss trends rather than derived from formal proofs.
+  15. Guidelines:
+    • The answer NA means that the paper does not include theoretical results.
+    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
+    • All assumptions should be clearly stated or referenced in the statement of any theorems.
+    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
+    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
+    • Theorems and Lemmas that the proof relies upon should be properly referenced.
+  16. 4.
+    Experimental result reproducibility
+  17.
+Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
+  18. Answer: [Yes]
+  19. Justification: All experimental setups, including model size, data budget, compute used, and training protocols, are fully described in Section. Code is opensourced.
+  20. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
+    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
+    • Depending on the contribution, reproducibility can be accomplished in various ways.
+For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model.
+In general.
+releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
+    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution.
+For example
+      1. (a)
+        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
+      2. (b)
+        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
+      3. (c)
+        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
+      4. (d)
+        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility.
+In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
+  21. 5.
+    Open access to data and code
+  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
+  23. Answer: [Yes]
+  24. Justification: Code and data is made public.
+  25. Guidelines:
+    • The answer NA means that paper does not include experiments requiring code.
+    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer.
+Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
+    • The instructions should contain the exact command and environment needed to run to reproduce the results.
+See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
+    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
+    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines.
+If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
+    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
+    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
+  26. 6.
+    Experimental setting/details
+  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
+  28. Answer: [Yes]
+  29. Justification: The paper explains the model hyperparameters and training setup.
+  30. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
+    • The full details can be provided either with the code, in appendix, or as supplemental material.
+  31. 7.
+    Experiment statistical significance
+  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
+  33. Answer: [Yes]
+  34. Justification: We showed the results with different amount of unique dataset sizes and compute.
+  35. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
+    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
+    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
+    • The assumptions made should be given (e.g., Normally distributed errors).
+    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
+    • It is OK to report 1-sigma error bars, but one should state it.
+The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
+    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
+    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
+  36. 8.
+    Experiments compute resources
+  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
+  38. Answer: [Yes]
+  39. Justification: The detailed information of the compute resources and configuration are provided in the supplementary materials.
+  40. Guidelines:
+    • The answer NA means that the paper does not include experiments.
+    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
+    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
+    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
+  41. 9.
+    Code of ethics
+  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
+  43. Answer: [Yes]
+  44. Justification: We discussed he broader impact of the work in the last section. of Appendix.
+  45. Guidelines:
+    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
+    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
+    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
+  46. 10.
+    Broader impacts
+  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
+  48. Answer: [Yes]
+  49. Justification: We discussed he broader impact of the work in the last section. of Appendix.
+  50. Guidelines:
+    • The answer NA means that there is no societal impact of the work performed.
+    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
+    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
+    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments.
+However, if there is a direct path to any negative applications, the authors should point it out.
+For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation.
+On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
+    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
+    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
+  51. 11.
+    Safeguards
+  52.
+Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
+  53. Answer: [N/A]
+  54. Justification: The models trained are small scale.
+  55. Guidelines:
+    • The answer NA means that the paper poses no such risks.
+    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
+    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
+    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
+  56. 12.
+    Licenses for existing assets
+  57.
+Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
+  58. Answer: [Yes]
+  59. Justification: The framework and datasets that we used to train and evaluate are open-sourced and correctly cited.
+  60. Guidelines:
+    • The answer NA means that the paper does not use existing assets.
+    • The authors should cite the original paper that produced the code package or dataset.
+    • The authors should state which version of the asset is used and, if possible, include a URL.
+    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
+    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
+    • If assets are released, the license, copyright information, and terms of use in the package should be provided.
+For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets.
+Their licensing guide can help determine the license of a dataset.
+    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
+    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
+  61. 13.
+    New assets
+  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
+  63. Answer: [Yes]
+  64. Justification: The provided code is well-documented.
+  65. Guidelines:
+    • The answer NA means that the paper does not release new assets.
+    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
+    • The paper should discuss whether and how consent was obtained from people whose asset is used.
+    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
+  66. 14.
+    Crowdsourcing and research with human subjects
+  67.
+Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
+  68. Answer: [N/A]
+  69. Justification: The paper does not involve any human subjects or crowdsourcing experiments.
+  70. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
+    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
+  71. 15.
+    Institutional review board (IRB) approvals or equivalent for research with human subjects
+  72.
+Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
+  73. Answer: [N/A]
+  74. Justification: The work does not involve human participants or subject studies.
+  75. Guidelines:
+    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
+    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research.
+If you obtained IRB approval, you should clearly state this in the paper.
+    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
+    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
+  76. 16.
+    Declaration of LLM usage
+  77.
+Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research?
+Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
+  78. Answer: [N/A]
+  79. Justification: No use of LLMs.
+  80. Guidelines:
+    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
+    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.